### v1

# Gas price shocks, Uncertainty and Price setting: Evidences from Italian Firms

Abstract
This paper examines how natural gas price shocks affect Italian firms’ pricing decisions and inflation expectations using quarterly survey data from the Bank of Italy’s Survey on Inflation and Growth Expectations (SIGE) spanning 1999Q4–2025Q2. We identify natural gas price shocks through a Bayesian VAR with sign and zero restrictions. Our findings reveal that these shocks are a primary driver of firms’ inflation expectations, particularly during the post-COVID period (2021–2023) when supply disruptions following Russia’s invasion of Ukraine generated unprecedented price pressures. We then estimate a larger BVAR incorporating firm-level price setting variables and macro aggregates, documenting that gas price shocks generate persistent increases in both firms’ current and expected prices, alongside elevated inflation uncertainty. We uncover substantial non-linearities using state-dependent local projections: under high uncertainty, firms successfully pass through cost increases to consumers, maintaining elevated prices; under low uncertainty, recessionary effects dominate, causing firms to reduce prices below baseline.
Keywords: Gas price shocks, Firms’ Price Setting, Inflation Uncertainty, Firms Expectations, Local Projections, Bayesian Vector Autoregression, Sign Restrictions
JEL Codes:E31, C20, D840, C11,Q41
1 Introduction
The period that follows the Covid-19 pandemic has seen an unprecedented rise in inflation in European economies. The years that followed the post pandemic recovery have been characterized by strong political tensions, dominated by events such as Russian invasion of Ukraine in the first part of 2022 which led to massive disruptions in natural gas supplies to European countries, raising the attention on inflationary role of supply shocks affecting this important commodity. Due to the lack of domestic energy sources, Italy is a major importer and deeply reliant on natural gas as one of its main power sources, making it particularly sensitive to supply disruptions that followed this dramatic geopolitical turning point. Italian firms have been particularly exposed to the unprecedented increase in the price of natural gas, which had a substantial impact on the price of electricity in the country. We show that, unsurprisingly, this has led to an unprecedented increase in firms’ uncertainty about the future level of prices.
Unlike the extensive literature on oil supply shocks, research examining the role of natural gas supply disruptions has emerged only recently. In this paper, we aim at identifying a shock to the nominal price of natural gas and at evaluating the impact of this shock on firms price setting behavior. We identify the shock by adopting the behavioral Bayesian VAR model proposed by Kilian and Zhou (2022), on which we impose a combination of sign and zero restrictions. Then, we evaluate the effect of this shock on firms’ pricing choices using generalized impulse response functions from a larger BVAR for the Italian economy. Additionally, we evaluate non linear effects in the transmission of the shock to firms’ prices by using state dependent local projections framework (Ramey and Zubairy, 2018; Falck, Hoffmann and Hürtgen, 2021). Our contribution to the literature lies in examining how natural gas price shocks affect firm-level outcomes using microdata. We base our analysis on quarterly survey data from the Bank of Italy’s Survey on Inflation and Growth Expectations (SIGE).
Related literature.
We draw from a vast literature about macroeconomic outcomes of energy related shocks. A large number of papers study the macroeconomic effects of oil shocks (Hamilton, 1983; Kilian, 2009; Caldara, Cavallo and Iacoviello, 2019; Conflitti and Luciani, 2019; Kilian and Zhou, 2022). Our work is closely related to the recent papers which study the effect of gas price and supply shocks, focusing on the recent inflationary surge in European economies. Alessandri and Gazzani (2025) identify a natural gas supply shock using daily news on European gas market as an instrument. Boeck and Zörner (2025) and López et al. (2025) estimate the pass through to inflation of gas price shocks. Adolfsen et al. (2024) use a BVAR framework to identify shocks driving the natural gas market in the EU. They document that the pass through is heterogeneous depending on the shock type. Güntner, Reif and Wolters (2024) study the effect of recent supply disruptions in natural gas supply in the German market. Casoli, Manera and Valenti (2024) study the interaction between oil/gas shocks and their effects on inflation in the Euro area. With respect to those authors, we focus on firm level expectations and price setting behavior.
Secondly, our work relates to the literature on uncertainty Bloom, Bond and Van Reenen (2007); Bloom (2009), especially with regard to that related to measuring uncertainty in agents’ expectations (Binder, 2017; Jurado, Ludvigson and Ng, 2015; Rossi and Sekhposyan, 2015; Manski, 2018) and to quantifying the effects of uncertainty on the macro economy (Bloom et al., 2018; Ascari and Haber, 2022; Georgarakos et al., 2024; Fasani et al., 2025).
The rest of the paper is organized as follows. In section 2 we describe the main features in the Survey of Inflation and Growth Expectaion (SIGE), which constitute our main data source of information for firm level microdata. We then describe the aggregate statistics obtained from the survey and that we use for conducting our empirical analysis, which is described in section 3. section 4 concludes.
2 Data: Survey on Inflation and Growth Expectations
This section illustrates the features of the Italian Survey on Inflation and Growth expectations (SIGE) [^1] Survey on Inflation and Growth Expectations, which constitutes our primary data source for firm expectations and price-setting decisions. As one of the longest-running firm-level expectation surveys in a G7 country, SIGE provides a rich source of information on both expectations formation and pricing behavior.
The survey is conducted by Bank of Italy at a quarterly frequency starting from the end of 1999. It provides a rotating panel collecting different kind of information from Italian firms. Among others, firms are asked to provide a point estimate for their year on year inflation expectations, and the expected change in their own prices over the course of the year. Furthermore, firms report the average change in their own realized price over the year, allowing a comparison between expectations and outcomes. Additionally, firms respond to a range of categorical question related to what are the main factors that will affect their own prices over the course of the next year. In this case, responses are on a scale giving information about both direction (downward or upward pressure) and intensity (ranging from strong to modest). In the following paragraphs, we describe the aggregate information that we extract from the survey in order to be used in our analysis.
Diffusion indexes
To determine which factors are most likely to impact firms’ price-setting behavior, we examine their responses to the categorical survey questions. Firms are asked about both the direction (positive, negative) and the intensity (strong, medium, modest) with which a specific factor is likely to affect their prices in the next year. The factors that are considered are the prices of raw materials and intermediate inputs, their inflation expectations for the next year, the prices of other competing firms, the trend in labor cost and aggregate demand and changes in the situation related to financing conditions. By using these responses, we build diffusion indexes using the approach of Pinto, Sarte and Sharp (2020). Those indexes, shown in Figure 1 together with 95% confidence bands, capture the average perceived intensity and direction of each factor across firms and are informative about movements in the distribution of firms’ responses. [^2] In order to buid the indexes, each response is weighted by its intensity on a -3 to +3 scale, where higher absolute values indicate stronger perceived pressure.
Figure 1: Diffusion indexes on factors affecting firms’ future prices. The figure reports diffusion indexes computed on the categorical questions in the SIGE survey about the factors that can affect a firm’s prices in the next year, togheter with 95% confidence bands. Source: Survey of Inflation and Growth Expectation. Sample: 1999:Q4–2025:Q2.
The observed patterns suggest the presence of a cost-push narrative during the European energy crisis that followed the post-pandemic recovery: raw material and intermediate input costs are cited as primary pricing drivers starting in 2021, followed by labor costs and demand pressures. Notably, inflation expectations themselves become a significant pricing factor only around 2022, suggesting that as firms observed widespread price increases, expectations began feeding directly into their own pricing decisions. Given the prominence of input costs—particularly energy-related raw materials—as key drivers of pricing behavior during this period, this motivates our focus on examining the effects of natural gas price shocks on firm-level outcomes.
Firms expectations and price setting
In this paragraph, we report the first moment of firms’ inflation expectations during the next year, firms expectations about the percentage of change in the price they will charge in the next year and the percentage of change of the price they charged during the current year. Starting from the third quarter of 2012, an important innovation was introduced in the survey. Prior to this change, all participating firms were informed about the current level of inflation, which was reported in the survey questionnaire. Since 2012Q3, the sample size has been increased and firms have been randomly assigned to two groups: one receiving updated information about current inflation, and the other receiving no such information. Due to the lenght of the sample size, the aggregate statistics we report are those related to firms which have been updated about the current level of prices. It is then not surprising that the average annual inflation expectations reported in Figure 2 is closely tied to realized inflation. [^3] This stress the importance of lack of information and inattention as key sources of bias in inflation expectations.
Figure 2: Average firms’ annual inflation expectations. The figure reports the first moment of firms’ year-on-year inflation expectations, as well as the realized HICP. Source: Survey of Inflation and Growth Expectation. Sample: 1999:Q4–2025:Q2.
However, it is possible to observe a detachment from actual realizations during periods characterized by huge levels of economic distress. During the Leheman crisis of 2008 and the sovereign debt crisis firms tend to over predict inflation, while in the aftermarth of the Covid pandemic they massively under predicted it.
Figure 3: Average change in firms’ annual expected and realized prices. The figure reports the first moment of firms’ year-on-year expectations on their own prices, as well as the realized change in the price they charge and HICP. Source: Survey of Inflation and Growth Expectation. Sample: 1999:Q4–2025:Q2.
Figure 3 shows the average firms’ expectations for their own prices in the upcoming year, as well as the annual change in the price they charge. In general, these measures are lower than actual inflation. Both of them increase after 2020. Expected price changes peaked in 2022, in correspondence of Russian invasion of Ukraine and the start of the European energy crisis, while realized prices adjusted more sluggishly. Notably, both expected and realized prices remained at an higher level with respect to realized HICP.
Inflation uncertainty
To measure firms’ inflation uncertainty, we adopt the approach proposed by Binder (2017), which exploits the well-documented tendency of survey respondents to provide round-number forecasts when they face greater uncertainty. This method yields an index that captures the proportion of likely uncertain firms in each survey wave. Since firms are informed about current price levels at the time of the interview, our index isolates doubts about future inflation, abstracting from any confusion about the current state of prices. Figure 4 reports inflation uncertainty index between the last quarter of 1999 and the second quarter of 2025.
Figure 4: Inflation Uncertainty Index. The figure reports the inflation uncertainty index for firms computed by adopting the framework of Binder (2017). Vertical red bars higlight moments of high uncertainty. Source: Survey of Inflation and Growth Expectation. Sample: 1999:Q4–2025:Q2.
Inflation uncertainty among informed firms is notably lower than that typically observed for consumers (Binder, 2017). During the early sample period, uncertainty remains low, reflecting stable price dynamics. The first significant spike occurs in 2008Q3 during the Global Financial Crisis (GFC), followed by elevated levels during the sovereign debt crisis. Subsequent peaks coincide with the Russian invasion of Crimea (2014Q3), Greece’s IMF default (2015Q2), and the peak in the Italian banking crisis (2016Q3). The fall of the Conte II government in 2020Q4 triggers a sharp increase in uncertainty, likely reflecting concerns about Italy’s access to European recovery funds during the COVID-19 pandemic. The most dramatic increase in uncertainty emerges with Russia’s invasion of Ukraine in 2022Q1, driven by severe energy supply disruptions and their inflationary consequences. Uncertainty peaks again in 2022Q4, as Italian inflation reaches its highest level since the introduction of the Euro, reflecting widespread concerns about the persistence and breadth of price pressures.
3 Empirical Analysis
3.1 Shock identification and impulse response functions
In order to identify a shock of nominal natural gas price, we propose the same framework as Kilian and Zhou (2022), but we apply it to the case of natural gas. We estimate a VAR with three variables: the real natural gas price, the log change in HICP, computed by excluding the price of natural gas, and firms year on year inflation expectations. The model is estimated at quarterly frequency from 1999Q4 to 2025Q2. Since the SIGE survey is conducted at the end of each quarter, all variables are measured at the same point in time—specifically, the final month of each quarter (March, June, September, December). This alignment preserves the within-month timing structure necessary for identification and ensures the proper alignment of the data.
Let $y_{t,m=3}=[\pi_{t,m=3}^{gas},\pi_{t,m=3}^{core},E_{t,m=3}\pi_{t+1}^{firms}]$ be a vector containing natural gas price $\pi_{t,m=3}^{gas}$, the log difference in HICP inflation excluding gas $\pi_{t,m=3}^{core}$ and year-on-year inflation expectations of firms $E_{t,m=3}\pi_{t+1}^{firms}$. The structural VAR model is given by:
$B_{0}y_{t,m=3}=\sum_{j=1}^{P}B_{j}y_{t-j,m=3}+w_{t,m=3}$ (1)
where $w_{t,m=3}$ is a vector of structural innovations. We can write the reduced form model as:
$y_{t,m=3}=\sum_{j=1}^{P}A_{j}y_{t-j,m=3}+u_{t,m=3}$ (2)
where $A_{j}=B_{0}^{-1}Bj$, j=1,2,,Pj=1,2,...,P, and the lag order $P$ is set to 4. We identify the nominal natural gas price shock by placing a combination of sign and zero restrictions on $B_{0}^{-1}$ by using the Bayesian approach of Arias, Rubio-Ramírez and Waggoner (2018) and adopting a uniform-Normal-inverse Wishart Prior. The nominal natural gas price shock is assumed to increase the real price of natural gas, as core inflation does to not respond in the same month as the price of natural gas. It increase headline inflation and inflation expectations of firms. The shock to core inflation reduces the real price of natural gas, since it does not respond immediately to an inflation shock, and increases core inflation and inflation expectations. Finally, a shock to firms’ inflation expectations has 0 effect on impact on real natural gas price and core inflation. The restrictions are reported in Equation 3
[…]
Figure 5 reports the impulse response functions to a one-standard-deviation natural gas price shock. The shock generates a sharp and persistent increase in the real natural gas price, with the effect gradually dissipating over approximately 10 quarters. The log change in headline inflation responds immediately, peaking on impact at around 15 basis points before returning to baseline within 5 quarters, consistent with the transitory nature of energy price shocks. In contrast, firms’ inflation expectations exhibit a more persistent response, with a positive response dying out in about 12 quarters.
Figure 5: Impulse response funcions of real natural gas price, headline inflation and firms’ inflation expectations to a natural gas price shock. The figure reports the impulse response functions of real natural gas price (left), headline inflation (center) and firms’ inflation expectations (right) to a natural gas price shock. Shaded areas represent 68% confidence bands. Sample: 1999:Q4–2025:Q2.
The impulse response functions show that natural gas price shocks are an important driver of firms’ inflation expectations. Figure 6 presents the historical decomposition of these expectations and of the log change in HICP excluding gas, isolating the contribution of each identified shock. Natural gas price shocks emerge as a primary driver of variation in firms’ annual inflation expectations throughout the sample period. These shocks predominantly exerted downward pressure on expectations, with a notable exception during the post-COVID-19 period (2021-2023), when they contributed substantially to the upward revision in firms’ inflation forecasts. This is unsurprising, given the high exposure of the Italian economy to natural gas supplies as its main energy source. Additionally, the appear to drive upward the log change in HICP during the same period.
Figure 6: Historical decomposition of firms’ year-on-year inflation expectations and log change in HICP excluding natural gas. The figure reports the historical decomposition of firms’ annual inflation expectations (top) and log change in HICP excluding natural gas price (bottom). Sample: 1999:Q4–2025:Q2.
3.2 Natural gas price shocks and firms’ pricing decisions
In order to evaluate how natural gas price innovations affect the pricing decisions of firms, we extract the identified structural shock from the first VAR. Then, we estimate a larger model, in which the shock is ordered as the first variable. Apart from natural gas price innovations, the model contains our estimated index of inflation uncertainty, the first moment of firms’ annual rate of change in their own prices and in their year-on-year expectations for their own prices, the annual HICP, unemployment rate, the short term interest rate for the Euro area and an industrial confidence index. The model is estimated by using the most advanced Bayesian techniques, adopting the hierarchical approach of Giannone, Lenza and Primiceri (2015) for the estimation of the hyperparameters and Lenza and Primiceri (2022) correction in order to account for the huge volatility of shocks during Covid. As all our variables are stationary, we estimate the model by shrinking the autoregressive term to 0. Our sample spans from the last quarter of 2002 until the second quarter of 2025. Due to the short sample and the large number of variables in the model, we estimate it with 2 lags. However, increasing the number of lags to 4 delivers nearly identical results.
Figure 7: Impulse response functions of aggregate macroeconomic variables to a natural gas price shock. The figure reports the impulse response functions of inflation uncertainty undex, firms’ expected and realized annual price change, annual HICP, unemployment rate, short term interest rate and industrial confindence index to a natural gas price shock. Shaded areas represent 68% confidence intervals. Sample: 2002:Q4–2025:Q2.
Figure 7 reports the generalized impulse response functions to a one standard deviation natural gas price innovation, obtained by imposing recursivity in the model. As a consequence of the shock, annual inflation increases by about 10%. The effect is persistent and vanishes after roughly six quarters. Firms react by raising both their own prices and their expected prices in the short run, pointing to a rapid pass-through of higher energy costs. These responses peak within the first two quarters and then gradually fade away as weaker demand conditions and tighter monetary policy curb further price increases. At the same time, inflation uncertainty rises, suggesting that the shock not only pushes up the level of prices but also complicates firms’ forecasting. The industrial confidence index shows a short-lived improvement in the immediate aftermath of the shock, possibly reflecting temporary gains for some firms able to pass on costs, but it turns negative after the fifth quarter as higher energy prices depress demand and profitability. Overall, the adjustment remains temporary: the cost-push nature of the shock dominates in the short run, but the inflationary pressures dissipate once the economy absorbs the shock.
3.3 State dependent local projections
In order to evaluate the presence of non-linearities in firms’ price setting behavior in response to a shock to the price of natural gas, we employ state dependent local projections (Ramey and Zubairy, 2018). We determine states of high and low firms’ inflation uncertainty using the index constructed using the firm level microdata from the SIGE survey by adopting the approach of Binder (2017). Following Falck, Hoffmann and Hürtgen (2021) we scale the uncertainy index for the average inflation expectations of firms, in order to account for high inflationary periods, and we smooth it by applying a 4 period backward looking weighted moving average filter. We adopt two different rules for determining states of high inflation uncertainty. According to the first rule, the transition between states of high and low probabilities of inflation uncertainty is governed a logistic function $Z(\hat{\Delta}_{t-1})\in[0,1]$ of the lagged index, defined as:
$Z(\hat{\Delta}_{t-1})=\frac{exp\left(\eta\frac{\hat{\Delta}_{t-1}-\mu}{\sigma_{\hat{\Delta}}}\right)}{1+exp\left(\eta\frac{\hat{\Delta}_{t-1}-\mu}{\sigma_{\hat{\Delta}}}\right)}$ (4)
where $\hat{\Delta}_{t}$ is the state variable, $\eta$ is a parameter that determines the steepness of the transition, and $\mu$ and $\sigma_{\hat{\Delta}}$ represent the median and standard deviation of the state variable, respectively. Following Falck, Hoffmann and Hürtgen (2021), we set $\eta=5$ and we compare probability regimes with the one obtained by adopting a simple threshold rule: the economy is in the high-uncertainty state $Z(\cdot)$ with probability 1 whenever $\hat{\Delta}_{t-1}>0$, and 0 otherwise.
Figure 8: Inflation uncertainty states. The figure displays the estimated probability of the high inflation uncertainty state, based on the measure by Binder (2017), and as defined by the smooth transition logistic function (light gray areas) and the threshold rule (dark gray areas). Sample: 2004:Q3–2025:Q2.
The transition probabilities are reported in Figure 8, together with the smoothed and scaled index of inflation uncertainty. The probability of being in the high uncertainty state increases in the aftermarth of the GFC and remains high during the sovereign debt crisis. It falls to 0 after 2019, and then it surges again in the aftermarth of the Covid-19 episode and during the European energy crisis, remaining high until the end of our sample. We then estimate the following set of state dependent local projections:
$x_{t+h}=$ $\bigg[\tilde{\beta}^{Z}_{x}(h)w^{gas}_{t}+\textit{controls}(x,Z)\bigg]Z(\hat{\Delta}_{t-1})$ $+\bigg[\tilde{\beta}^{\bar{Z}}_{x}(h)w^{gas}_{t}+\textit{controls}(x,\bar{Z})\bigg]\bar{Z}(\hat{\Delta}_{t-1})+\epsilon_{t+h}$ (5)
Where $\bar{Z}\equiv 1-Z(\hat{\Delta}_{t-1})$. The estimated impulse response function are reported in Figure 9. The top row shows the impulse responses obtained by adopting the simple threshold state transition rule, while the responses at the bottom are obtained by adopting the smooth transition rule based on the logistic function of Equation 4. Strong nonlinear effects emerge in the transmission of natural gas price shocks to firms’ price-setting decisions. Under the high uncertainty scenario, firms are confident that they will be able to pass the increased costs to their buyers. They increase their own prices persistently and raise their expectations for next year’s prices. However, this confidence appears to be short-lived. As the recession hits, price increases stop. Under the low uncertainty scenario, by contrast, firms anticipate the recessionary impact of the shock. They revise their prices downward and adjust their expectations accordingly.
Threshold rule
Smooth transition rule
Figure 9: Firms expected and realized annual price change state-dependent response to a natural gas price shock. Estimation follows the lag-augmentation method of Montiel Olea and Plagborg‐Møller (2021), with 4 lags. The responses in the top row are estimated using a clear threshold rule. The responses at the bottom row are estimated using the smooth transition probability rule. Shaded bands represent 68% confidence intervals based on Newey-West standard errors. Sample: 2004Q3–2025Q2.
4 Conclusions
This paper provides new evidence on how natural gas price shocks affect firms’ pricing decisions and inflation expectations in Italy, a country heavily reliant on natural gas as its primary energy source. Using quarterly survey data from the Bank of Italy’s SIGE spanning over two decades, we identify structural natural gas price shocks through a Bayesian VAR framework with sign and zero restrictions. Our analysis reveals three main findings. First, natural gas price shocks are a significant driver of firms’ inflation expectations throughout the sample period, with particularly strong effects during the post-COVID energy crisis of 2021-2023. Historical decomposition shows that these shocks predominantly exerted downward pressure on expectations during normal times but contributed substantially to the surge in inflation expectations following Russia’s invasion of Ukraine. Second, firms respond to natural gas price innovations by adjusting both their current and expected own prices, with inflation uncertainty rising following the shock. Both aggregate inflation and firms’ own prices exhibit similar dynamics, with effects dissipating within 5-6 quarters, consistent with the transitory nature of energy price shocks documented in the literature. Inflation uncertainty also increases in response to the shock, suggesting that energy price disruptions affect not only the level of prices but also firms’ confidence in forecasting future inflation. Third, and most importantly, we document substantial non-linearities in firms’ responses that depend on the pre-existing level of inflation uncertainty. When uncertainty is high, firms successfully pass through cost increases to consumers, maintaining elevated prices throughout the adjustment period. In contrast, when uncertainty is low, the recessionary effects of the shock dominate, causing firms to reduce prices below baseline as weak demand conditions prevent cost pass-through.