### v3

# Gas supply shocks, Uncertainty and Price setting: Evidences from Italian Firms

Abstract
This paper examines how natural gas supply shocks affect Italian firms’ pricing decisions and inflation expectations using quarterly survey data from the Bank of Italy’s Survey on Inflation and Growth Expectations (SIGE). We identify natural gas supply shocks through an external IV-VAR approach exploiting likely unexpected news about interruption to gas supplies to Europe. Our findings show that although gas supply shocks do not have huge effects on gas quantity and only modest effect on gas inventories, they are quickly transmitted to spot electricity prices with persistent effects. We then estimate a proxy internalizing BVAR incorporating firm-level variables from SIGE, documenting that gas supply shocks raise firms’ current and expected prices as well as inflation uncertainty. Finally, we uncover substantial nonlinearities using state-dependent local projections: under high inflation uncertainty, firms successfully pass cost increases on to consumers, sustaining elevated prices; under low uncertainty, recessionary effects dominate, leading firms to cut prices below baseline.
Keywords: Gas supply shocks, Firms’ Price Setting, Inflation Uncertainty, Firms Expectations, Local Projections, Bayesian Vector Autoregression, Instrumental Variables
JEL Codes:E31, C20, D840, C11,Q41
1 Introduction
The period that follows the Covid-19 pandemic has seen an unprecedented rise in inflation in European economies. The post pandemic recovery has been characterized by strong political tensions, dominated by events such as Russian invasion of Ukraine in the first part of 2022 which led to massive disruptions in natural gas supplies to European countries, raising the attention on inflationary role of supply shocks affecting this important commodity. Due to its limited domestic energy resources, Italy is a major importer and remains highly dependent on natural gas as a primary source of power, making it particularly sensitive to the supply disruptions that followed this dramatic geopolitical turning point. Italian firms have been particularly exposed to the unprecedented increase in the price of natural gas, which had a substantial impact on the price of electricity in the country. We show that, unsurprisingly, the post-pandemic recovery period has been characterized by an unprecedented increase in firms’ uncertainty about the future level of prices.
Unlike the extensive literature on oil supply shocks, research examining the role of natural gas supply disruptions has emerged only recently. In this paper, we aim at identifying a shock to the supply of natural gas and at evaluating the impact of this shock on firms price setting behavior. We identify the shock following an external IV-VAR approach following kanzig2021macroeconomic. We construct the instrument by exploiting episodes that carried likely exogenous news about supply distruptions affecting Europe, and in particular Italy. Our identification strategy is hence close in the spirit to the one followed by alessandri2025natural, which instead adopt an internal instrument and consider a larger number of surprises. Then, we evaluate the effect of this shock on firms’ pricing behavior and expectations using the generalized impulse response functions from a proxy internalizing VAR, using the shock identified in the first VAR as an external instrument. Additionally, we evaluate non linear effects in the transmission of the shock to firms’ prices by using state dependent local projections framework (ramey2018government; F2021). Our contribution to the literature lies in examining how natural gas supply shocks affect firm-level outcomes and the pass-through using firm level micro data. We base our analysis on quarterly data from the Bank of Italy’s Survey on Inflation and Growth Expectations (SIGE).
Related literature.
We draw from a vast literature about macroeconomic outcomes of energy related shocks. A large number of papers study the macroeconomic effects of oil shocks (hamilton1983oil; kilian2009not; caldara2019oil; conflitti2019oil; kilian2022oil). Our work is closely related to the recent papers which study the effect of gas supply shocks, focusing on the recent inflationary surge in European economies. alessandri2025natural identify a natural gas supply shock using daily news on European gas market as an instrument. boeck2025natural and lopez2025pass estimate the pass through to inflation of gas price shocks. adolfsen2024gas use a BVAR framework to identify shocks driving the natural gas market in the EU. They document that the pass through is heterogeneous depending on the shock type. guntner2024sudden study the effect of recent supply disruptions in natural gas supply in the German market. casoli2024energy study the interaction between oil/gas shocks and their effects on inflation in the Euro area. With respect to those authors, we focus on firm level expectations and price setting behavior.
Secondly, our work relates to the literature on uncertainty bloom2007uncertainty; Bloom2009, especially with regard to that related to measuring uncertainty in agents’ expectations (binder2017measuring; Jurado2015; RossiTatevik; Manski2018) and to quantifying the effects of uncertainty on the macro economy (Bloom2018; ascari2022non; GGCK2024; fasani2025belief).
The rest of the paper is organized as follows. In section 2 we describe the main features in the Survey of Inflation and Growth Expectation (SIGE), which constitute our main data source of information for firm level micro data. We then describe the aggregate statistics obtained from the survey and that we use for conducting our empirical analysis, which is described in section 3. section 4 concludes.
2 Data: Survey on Inflation and Growth Expectations
This section illustrates the features of the Italian Survey on Inflation and Growth expectations (SIGE) [^1] Survey on Inflation and Growth Expectations, which constitutes our primary data source for firm expectations and price-setting decisions. As one of the longest-running firm-level expectation surveys in a G7 country, SIGE provides a rich source of information on both expectations formation and pricing behavior.
The survey is conducted by Bank of Italy at a quarterly frequency starting from the end of 1999. It provides a rotating panel collecting different kind of information from Italian firms. Among others, firms are asked to provide a point estimate for their year on year inflation expectations, and the expected change in their own prices over the course of the year. Furthermore, firms report the average change in their own realized price over the year, allowing a comparison between expectations and outcomes. Additionally, firms respond to a range of categorical question related to what are the main factors that will affect their own prices over the course of the next year. In this case, responses are on a scale giving information about both direction (downward or upward pressure) and intensity (ranging from strong to modest). In the following paragraphs, we describe the aggregate information that we extract from the survey in order to be used in our analysis.
Diffusion indexes
To determine which factors are most likely to impact firms’ price-setting behavior, we examine their responses to the categorical survey questions. Firms are asked about both the direction (positive, negative) and the intensity (strong, medium, modest) with which a specific factor is likely to affect their prices in the next year. The factors that are considered are the prices of raw materials and intermediate inputs, their inflation expectations for the next year, the prices of other competing firms, the trend in labor cost and aggregate demand and changes in the situation related to financing conditions. By using these responses, we build diffusion indexes using the approach of PintoSarteSharp2020. Those indexes, shown in Figure 1 together with 95% confidence bands, capture the average perceived intensity and direction of each factor across firms and are informative about movements in the distribution of firms’ responses. [^2] In order to build the indexes, each response is weighted by its intensity on a -3 to +3 scale, where higher absolute values indicate stronger perceived pressure.
Figure 1: Diffusion indexes on factors affecting firms’ future prices. The figure reports diffusion indexes computed on the categorical questions in the SIGE survey about the factors that can affect a firm’s prices in the next year, together with 95% confidence bands. Source: Survey of Inflation and Growth Expectation. Source: Bank of Italy, Survey on inflation and growth expectations. Sample: 1999:Q4–2025:Q2.
The observed patterns suggest the presence of a cost-push narrative during the European energy crisis that followed the post-pandemic recovery: raw material and intermediate input costs are cited as primary pricing drivers starting in 2021, followed by labor costs and demand pressures. Notably, inflation expectations themselves become a significant pricing factor only around 2022, suggesting that as firms observed widespread price increases, expectations began feeding directly into their own pricing decisions. Given the prominence of input costs—particularly energy-related raw materials—as key drivers of pricing behavior during this period, this motivates our focus on examining the effects of natural gas supply shocks on firm-level outcomes.
Firms expectations and price setting
In this paragraph, we report firms’ one-year-ahead inflation expectations, their expected price changes over the same horizon, and their actual price changes during the current year. Starting from the third quarter of 2012, an important innovation was introduced in the survey. Prior to this change, all participating firms were informed about the current level of inflation, which was reported in the survey questionnaire. Since 2012Q3, the sample size has been increased and firms have been randomly assigned to two groups: one receiving updated information about current inflation, and the other receiving no such information. Due to the length of the sample size, the aggregate statistics we report are those related to firms which have been updated about the current level of prices. It is then not surprising that the average annual inflation expectations reported in Figure 2 is closely tied to realized inflation. [^3] This stress the importance of informational frictions and inattention as key sources of bias in inflation expectations.
Figure 2: Average firms’ annual inflation expectations. The figure reports the first moment of firms’ year-on-year inflation expectations, as well as the realized HICP. Source: Survey of Inflation and Growth Expectation. Source: Bank of Italy, Survey on inflation and growth expectations, EUROSTAT. Sample: 1999:Q4–2025:Q2.
However, it is possible to observe a detachment from actual realizations during periods characterized by huge levels of economic distress. During the Lehman crisis of 2008 and the sovereign debt crisis firms tend to over predict inflation, while in the aftermath of the COVID pandemic they massively under predicted it.
Figure 3: Average change in firms’ annual expected and realized prices. The figure reports the first moment of firms’ year-on-year expectations on their own prices, as well as the annual realized change in the price they charge and realized HICP. Source: Bank of Italy, Survey on inflation and growth expectations, EUROSTAT. Sample: 1999:Q4–2025:Q2.
Figure 3 shows the average firms’ expectations for their own prices in the upcoming year, as well as the annual change in the price they charge. In general, these measures are lower than actual inflation. Both of them increase after 2020. Expected price changes peaked in 2022, in correspondence of Russian invasion of Ukraine and the start of the European energy crisis, while realized prices adjusted more sluggishly. Notably, both expected and realized prices remained at an higher level with respect to realized HICP.
Inflation uncertainty
To measure firms’ inflation uncertainty, we adopt the approach proposed by binder2017measuring, which exploits the well-documented tendency of survey respondents to provide round-number forecasts when they face greater uncertainty. This method yields an index that captures the proportion of likely uncertain respondents in each survey wave. Differently from her case, our respondents can provide decimal answers. Therefore, we need to adapt her framework to our setting by adjusting the likelihood function accordingly. We assume that there are two type of respondents. TypehType-h respondents are characterized by an high degree of uncertainty in their subjective probability distribution, while typeltype-l respondents by a low degree of uncertainty. TypehType-h respondents are assumed to give point estimates for their expected level of prices coming from the set $S_{h}=\{0,5,10,15,20,25\}$, while low uncertainty firms are allowed to answer with decimal numbers on the support $\mathbb{R}^{+}$. If an answer $R_{it}\notin S_{h}$ we can surely classify it as typeltype-l. Conversely, if $R_{it}\in S_{h}$ we need to estimate the probability that respondent $i$ is uncertain. We assume that the cross sectional density is the mixture of two distributions. For typehtype-h respondents, coming from $S_{h}$ the probability mass function is obtained by integrating the latent density over rounding bins: $\phi_{h}(s)=P(R_{i,t}=s|type-h)=\int_{s-\delta/2}^{s+\delta/2}p_{h}(x)dx$, with $j\in S_{h}$ and $\delta$ being the bin width. [^4] We set $\delta=0.1$. For typeltype-l respondents the probability density is simply $f_{l}(x)=p_{l}(y)$ with $y\in\mathbb{R}^{+}$ and $\phi_{l}(s)=\int_{s-\delta/2}^{s+\delta/2}p_{l}(x)dx$ for rounded observations. The cross sectional log likelihood will then be given by:
$l(\theta_{t})=\sum_{s=S_{h}}n_{h}(s)log[(\lambda_{t}\phi_{h}(s)+(1-\lambda_{t})\phi_{l}(s)]+\sum_{i=1}^{N_{c}}log[(1-\lambda_{t})f_{l}(y_{c,i};\mu_{l},\sigma_{l})]$ (1)
where $f_{l}$ is assumed to be normal, $n_{h}$ is the number of firms which responses belong to $S_{h}$, $N_{c}$ is the number of non rounded responses and $\theta=(\phi_{h},\lambda_{t},\mu_{l},\sigma_{l})$ is a vector of parameters . The mixture weight $\lambda_{t}$ is the portion of uncertain respondents.555binder2017measuring includes non respondents and measures the proportion of uncertain households as $U_{t}=\frac{\lambda_{t}N_{t}+\gamma N_{t}^{DK}}{N_{t}+\gamma N_{t}^{DK}}$ where $N_{t}$ is the total number of respondents, $N_{t}^{DK}$ is the number of respondents which answer is ”don’t know” and $\gamma$ is a parameter which measures their weight. As in the SIGE survey the question about expected inflation is mandatory, we do not have DKtypeDK-type respondents in our sample and hence we measure the proportion of uncertain firms by simply taking $\lambda_{t}$. The vector of parameters $\theta_{t}$ is estimated via maximum likelihood. Figure 4 reports the inflation uncertainty index between the last quarter of 1999 and the second quarter of 2025. Notably, since firms are informed about current price levels at the time of the interview, our index isolates doubts about future inflation, abstracting from any confusion about the current state of prices.
Figure 4: Inflation Uncertainty Index. The figure reports the inflation uncertainty index for firms computed by adopting the framework of binder2017measuring. Vertical red bars highlight moments of high uncertainty. Source: Bank of Italy, Survey on inflation and growth expectations. Sample: 1999:Q4–2025:Q2.
Inflation uncertainty among informed firms is notably lower than that typically observed for consumers (binder2017measuring). During the early sample period, uncertainty remains low, reflecting stable price dynamics. The first major spike occurs in 2008Q3 during the Global Financial Crisis, followed by persistently high levels throughout the European sovereign debt crisis. Subsequent surges are observed around the Russian invasion of Crimea (2014Q3), Greece’s IMF default (2015Q2), and during the Italian banking crisis (2016Q3). The fall of the Conte II government in 2020Q4 coincides with a sharp rise in uncertainty, likely reflecting concerns over Italy’s access to European recovery funds amid the COVID-19 pandemic. A major increase appears in 2022Q1 with Russia’s invasion of Ukraine, driven by severe energy supply disruptions and their inflationary effects. Uncertainty rises again in 2022Q4, as Italian inflation reaches its highest level since the adoption of the euro, underscoring widespread concerns about the persistence and spread of price pressures.
3 Empirical Analysis
3.1 Identification through external instrument
Instrument construction
In order to identify a supply shock to natural gas, we follow the approach of kanzig2023unequal (kanzig2021macroeconomic, kanzig2023unequal). As alessandri2025natural, we exploit news extracted from Refinitiv. Differently from them, we focus on a smaller number of news from Reuters about interruptions of gas supplies to Europe which are highly likely to be exogenous to the economy and unexpected by the markets. We include episodes of unexpected interruptions caused by extreme weather event, like the cold snap that hit Russia in 2012 forcing it to reduce the export of natural gas to satisfy its domestic demand, and interruptions in the supply caused by terrorist attacks, protests and unplanned maintenance in Libya and Algeria. Furthermore, we include incidents like the explosion that hit the Baumgarten natural gas facility in Austria in 2017 causing the interruption of the flow of gas to the Italian Hub of Tarvisio and the fire in the Freeport liquefied natural gas facility in US which reduced the supply of LNG to Europe. Additionally, for the 2022 period we include some announcements from Russia which have been perceived as a surprise from the market and which contain news about future reduction in the supply of natural gas. As an example, the announcement by Putin that Russia was going to stop the deliveries of natural gas, unless the payments were made in robles in March 2022 concretely reduced the amount of purchasable commodity for countries which could not or did not want to comply with the request, and has been followed by a positive movement of the Dutch TTF (alessandri2025natural). We avoid to include the sabotage of the Nord Stream 2 pipeline since the latter never actually delivered natural gas to Europe, thus the event did not concretely change the amount of commodity supplied to European countries. We also exclude the strikes in gas production facilities in Norway during 2022, since these episodes were motivated by the request of Norvegian workers to increase wages in response to inflationary pressures. Furthermore, to reduce the contamination of the instrument with geopolitical shocks we do not include the Russian invasion of Ukraine in February 2024 even though it has been followed by a massive surge in the Dutch TTF.
Table 1: Natural gas surprises
Event date Event description
1 January 2006 Russia stops flows to Ukraine due a payment dispute, blocking gas flows to Italy and Austria
22 February 2011 Due to protests and civil war, the Greenstream pipeline stopped delivering gas from Libya to Italy
31 January 2012 An unexpected cold snap forces Russia to reduce the supply of gas to Europe to satisfy its domestic demand
17 January 2013 Supplies from Algeria to Italy fall because of a terrorist attack in Almenas Gas Field
4 March 2013 Clashes with militias in Mellitha complex in Libya completely stops the gas flows to Italy.
30 September 2013 Protests in Libya cause the shut down of a gas pumping station, halting the gas flows to Italy
23 January 2014 Gas flows from Algeria to Italy stopped due to quality issues. Received gas had to be sent back
29 April 2015 Libya shuts down a gas field in the east due to protests
13 December 2017 A blast in Austrian hub of Baumgarten interrupts gas transit to Tarvisio hub in Italy
05 February 2022 An unplanned maintenance at Mellitah complex considerably reduces gas flows from Libya to Italy
23 March 2022 Putin asks European countries to pay gas deliveries in Rubles otherwise gas flows will stop
8 June 2022 Fire in the Freeport LNG facility in the US
14 June 2022 NS1 operate at limited capacity due to turbines stuck in Canada
31 August 2022 Flows of gas from NS1 shut down permanently
2 October 2022 Russia halts gas deliveries to Italy trough the TAG pipeline
  • Notes: The table reports the natural gas surprises we use to construct the instrument.
We end up isolating 15 surprises which are reported in Table 1. We construct a daily instrument which takes value of 1 in correspondence of any of the surprises. Then, we aggregate the instrument at the monthly level.
IV-VAR specification
We specify a monthly VAR including six variables. We take the day ahead Dutch TTF as a measure of the spot price of natural gas. Then, in the spirit of adolfsen2024gas, we include the Italian gas quantities, gas inventories and industrial production. We also control for the world industrial production and for the Prezzo Unico Nazionale (PUN) as a measure of the Italian spot electricity price. [^6] Data about gas quantities and gas inventories are provided by the Italian Ministero dell’ambiente e della sicurezza energetica. Gas quantities are defined as national production, plus imports minus exports. Gas inventories are given by the cumulative sum of the variation in inventories, given a base value of 7000 millions of standard cubic meters. Both gas quantities and gas inventories are seasonally adjusted by using the X-13 ARIMA seasonal adjustment method. Data about the PUN are provided by Gestore Mercati Energetici. World industrial production is provided by BH2019. Monitoring the response of the PUN to a gas supply shock is informative because, in a uniform-price electricity market with gas-fired units often setting the marginal cost, like the Italian one, changes in the TTF day-ahead price feed into generators’ bids and hence the Italian spot price with short lags; the PUN therefore reflects downstream pass-through from gas to electricity price. Controlling for world industrial production purges the gas-price equation of global demand pressure, sharpening identification of a gas supply shock. The VAR is estimated in log levels with a constant from 2004M1 until 2025M6. The lag order is set at 12. We assess instrument relevance by performing an F-test in the regression of the TTF residuals from the VAR on the instrument, following olea2021inference. An F-statistic above 10 indicates a strong instrument. In our case, the F-statistic equals 13.6, confirming high instrument strength.
Figure 5: Gas supply shock. The figure shows the natural gas supply shock obtained from the IV-VAR (left) and the corresponding correlogram (right). Sample: 2005:M1–2025:m6.
We also conduct the F-test allowing for heteroskedasticity, which delivers an F-statistic of 15.18.777The instrument explains around 4.9% of the TTF residual. Figure 5 reports the structural shock extracted from the IV-VAR and the corresponding correlogram.
Shock exogeneity
We test the exogeneity of our shock with respect to oil and energy-related shocks identified in the literature using a Wald test of joint significance. Specifically, we regress our shock measure on the contemporaneous and 12 lagged values of each structural innovation considered, as well as on 12 of its own lags. We consider the oil price news shock of Kanzig2020, the carbon price shock of kanzig2023unequal, the oil demand shocks of CALDARA20191 and BH2019, the oil supply,inventories and demand shock of BH2019, the oil price expectation shock of Baumeister2023, and the oil supply shocks of CALDARA20191, kilian2009not and kilian2008exogenous. Results are reported in Table 2. The Wald tests do not reject the null of no contamination for any proxy, supporting their exogeneity. Only the BH2019 oil inventories shock approaches significance at the 10% level.
Table 2: Wald test results.
Innovation Sample Wald statistic P-value
Kanzig2020 Oil price news 2000Q4 - 2024Q2 7.8868 0.7939
kanzig2023unequal Carbon price 2000Q4 - 2019Q4 16.0719 0.188
CALDARA20191 Oil demand 2000Q4 - 2015Q4 12.0753 0.4396
CALDARA20191 Oil supply 2000Q4 - 2015Q4 10.8998 0.5375
BH2019 Oil supply 2000Q4 - 2025Q2 9.2402 0.6823
BH2019 Oil consumption demand 2000Q4 - 2025Q2 11.9281 0.4515
BH2019 Oil inventories 2000Q4 - 2025Q2 20.1281 0.0647
Baumeister2023 Oil price expectation 2000Q4 - 2023Q1 6.1135 0.9102
kilian2009not Oil supply 2000Q4 - 2009Q4 5.7186 0.9296
kilian2008exogenous Oil Supply 2000Q4 - 2016Q4 8.2033 0.7690
  • Notes: The table reports Wald test statistics for the joint significance of coefficients. The null hypothesis is that all coefficients are zero, indicating no contamination.
Energy Market Responses to a Gas Supply Shock
Figure 6 shows the impulse response functions from the VAR estimated by using the external instrument approach. As in kanzig2021macroeconomic, we can interpret the impulse response functions as elasticities, as all the variables in the VAR are in logs. We normalize the shock to increase the gas price, measured by the TTF, by 1%. We can see that price variables immediately react to the shock. The TTF responds positively and the response reverts to 0 in a bit more than an year. The increase in the price of natural gas is then transmitted to the national electricity price, measured by the PUN, which increases strongly, peaking within a few months and then slowly reverting. This confirms pass-through from gas to electricity: gas-fired plants set the marginal price, so the electricity market mirrors gas dynamics with a short lag. About half of the increase in the gas price is transmitted to the electricity price on impact. After a few months, gas inventories decrease of about 0.1%. Higher gas prices coincide with withdrawals from storage before restocking once prices ease. Gas quantity initially do not respond to the shock. After 10 months, we observe a small drop in gas quantity followed by a gradual recovery. Industrial production is not affected by the shock, consistently with its transitory nature.
Gas supply shock
[Uncaptioned image] Figure 6: Impulse response functions of energy market variables to a natural gas supply shock. The figure reports the impulse response functions of natural gas price, gas inventories, gas quantity, industrial production, PUN and world industrial production to a natural gas supply shock. Shaded areas represent 68% and 90% confidence intervals, computed using a moving block bootstrap procedure. Source: Ministero dell’ambiente e della sicurezza energetica, Gestore Mercati Energetici, ISTAT, Refinitiv Workspace, BH2019. Sample: 2002:Q4–2025:Q2.
3.2 Natural gas supply shocks and firms’ behavior
Bayesian VAR
In order to evaluate how natural gas supply innovations affect the pricing decisions of firms, we estimate model shown in Section 3.1 and we extract the identified structural shock. Then, we estimate a proxy-internalizing VAR, placing the shock of interest as the first variable in the system and imposing recursivity (BRULU2022). [^8] According to PMW2021, the advantage of this approach over the equivalent SVAR-IV method of SW2018, besides its simplicity, is that the estimated impulse responses are valid even if the shock of interest is non-invertible. Apart from our estimated innovation, the model contains our index of inflation uncertainty, the average firms’ annual rate of change in their own prices and the average expected change in their prices over the next year, the annual rate of change in their assessment about order-book levels and about the stock of finished products and their inflation expectations. Firms expectations from SIGE are measured at quarterly frequency, at the final month of each quarter. In order to preserve the proper time alignment, we follow AB2023 and we take the end of the quarter values of the gas supply shock. Our sample spans from 2005Q4 until 2025Q2.
Gas supply shock
[Uncaptioned image] Figure 7: Impulse response functions of firms’ related variables to a natural gas supply shock. The figure reports the impulse response functions of inflation uncertainty index, firms’ expected and realized annual price change, firms’ assessed order-book levels and stock of finished products and annual inflation expectations to a natural gas supply shock. Shaded areas represent 68% confidence intervals. Source: Bank of Italy, Survey on inflation and growth expectations, EUROSTAT, European Commission - ECFIN. Sample: 2005:Q4–2025:Q2.
Due to the limitations of our sample size, we resort to Bayesian methods for recovering the impulse response functions. We adopt the hierarchical approach of giannone2015prior for the optimization of the hyperparameters and the lenza2022estimate correction, in order to account for the huge volatility of shocks during Covid. As all our variables are stationary, we estimate the model by shrinking the autoregressive term to 0. We set the number of lags to 4. Figure 7 reports the generalized impulse response functions to a shock that increases the TTF by 1%. All dependent variables in this specification are expressed in annual percentage changes. The results are consistent with a cost-push disturbance. On impact, the annual rate of change in firms’ prices increases by 5% and the effect fades in a bit more than one year. Expectations about future prices display a similar but shorter-lived pattern. The year on year change jumps higher with respect to price realizations on impact, and reverts to 0 in half an year. Inflation uncertainty increases sharply for about one year, suggesting that the shock not only pushes up the level of prices but also complicates firms’ forecasting. Inflation expectations move in the same direction and with a similar scale as price expectations, but the increase in their annual growth rate is as persistent as the one for firms’ realized prices.
State dependent local projections
adolfsen2024gas argue that the pass-trough of a natural gas supply shock in the Euro Area is characterized by a non linearity in the level of tightness in the labor market, with effects that are considerably magnified when the economy operates near full capacity. We show that the pass-through of natural gas supply shocks is also characterized by substantial non-linearities related firms’ ability to accurately forecast inflation. When inflation uncertainty increases, it becomes more difficult for firms to form an expectation about the future level of prices. This affects the way in which they respond to the shock of the price of natural gas. Under high uncertainty scenarios, the pass through is larger and firms persistently increase their prices. When instead uncertainty about future inflation is low they reduce their prices in order to preserve their market shares, anticipating the recessionary effects of the shock. We illustrate this by adopting state dependent local projections (ramey2018government). We determine states of high and low firms’ inflation uncertainty using the index constructed using the firm level micro data from the SIGE survey by adopting the approach of binder2017measuring. Following F2021 we scale the uncertainty index for the average inflation expectations of firms, in order to account for high inflationary periods, and we smooth it by applying a 4 period backward looking weighted moving average filter. We assume that the transition between states of high and low probabilities of inflation uncertainty is governed a logistic function $Z(\hat{\Delta}_{t-1})\in[0,1]$ of the lagged index, defined as:
$Z(\hat{\Delta}_{t-1})=\frac{exp\left(\eta\frac{\hat{\Delta}_{t-1}-\mu}{\sigma_{\hat{\Delta}}}\right)}{1+exp\left(\eta\frac{\hat{\Delta}_{t-1}-\mu}{\sigma_{\hat{\Delta}}}\right)}$ (2)
where $\hat{\Delta}_{t}$ is the state variable, $\eta$ is a parameter that determines the steepness of the transition, and $\mu$ and $\sigma_{\hat{\Delta}}$ represent the median and standard deviation of the state variable, respectively. Following F2021, we set $\eta=5$. [^9] Results are consistent across different values for parameter $\eta$. In order to account for the non linearities in unemployment shown by adolfsen2024gas, we do the same for the unemployment rate, determining the probabilities of having a tight labor market.
Figure 8: Inflation uncertainty and labor market states. Top panel: estimated probability of high inflation uncertainty following binder2017measuring, computed via smooth transition logistic function (shaded areas). Bottom panel: estimated probability of tight labor market based on unemployment, computed via smooth transition logistic function (shaded areas). Blue lines represent the underlying state variables. Source: Bank of Italy, Survey on inflation and growth expectations, EUROSTAT. Sample: 2007:Q1–2025:Q2.
The transition probabilities are reported in Figure 8, together with the smoothed and scaled index of inflation uncertainty and the unemployment rate. The probability of being in the high uncertainty state increases in the aftermath of the GFC and remains high during the sovereign debt crisis. It falls to 0 after 2019, and then it surges again in the aftermarth of the Covid-19 episode and during the European energy crisis, remaining high until the end of our sample. Instead, the probabilities of having a tight labor market are high until 2011. Then they remain at 0 until 2021, with the exception of a single episode, coinciding with the onset of the COVID pandemic in 2020. We then estimate the following set of state dependent local projections:
$x_{t+h}=$ $\bigg[\tilde{\beta}^{Z}_{x}(h)w^{gas}_{t}+\textit{controls}(x,Z)\bigg]Z(\hat{\Delta}_{t-1})$ $+\bigg[\tilde{\beta}^{\bar{Z}}_{x}(h)w^{gas}_{t}+\textit{controls}(x,\bar{Z})\bigg]\bar{Z}(\hat{\Delta}_{t-1})+\gamma_{D}D_{t}+\epsilon_{t+h}$ (3)
Where $Z(\hat{\Delta}_{t-1})$ indicate the probability of being in the state of high inflation uncertainty, $\bar{Z}\equiv 1-Z(\hat{\Delta}_{t-1})$, $\textit{controls}(x,\bar{S})$ is a set of control variables including the interaction between $w^{gas}_{t}$ and the probability of being in the tight labor market state $S(\hat{\Delta}_{t-1})$ and lags of the shock and of the dependent variable $x_{t}$. [^10] In our baseline specification, we include two lags of the dependent variable and two lags of the shock to preserve degrees of freedom. Results are robust to specifications with four lags of both variables. $D_{t}$ is a dummy variable that assumes value of 1 in correspondence of 2020Q2 and then decays exponentially over the subsequent quarters. The estimated impulse response function are reported in Figure 9.
Gas supply shock
Figure 9: Firms annual price change state-dependent response to a gas supply shock. The figure shows the state dependent response of firms’ prices to a gas supply shock. Estimation follows the lag-augmentation method of MOPM2021. The number of lags of the dependent variable, the control variables and the shock in each state are determined by minimizing the AIC. Shaded bands represent 68% confidence intervals based on Newey–West standard errors. Source: Bank of Italy, Survey on inflation and growth expectations, EUROSTAT. Sample: 2007Q1–2025Q2.
When inflation uncertainty is high, firms respond to a natural gas supply shock by persistently raising their prices. This behavior reflects precautionary pricing motives: firms face higher perceived risks about future input costs and the duration of the energy shock, and therefore widen their mark-ups to protect margins against potential future cost increases. The widespread and salient nature of energy costs also facilitates price coordination across firms, reducing competitive pressure and allowing a stronger pass-through of cost shocks. Under low uncertainty instead, firms’ prices fall in response to the shock to the gas supply. The negative effect of the shock on aggregate demand is clear. Firms anticipate the recession and do not increase their prices in order to preserve their market share.
4 Conclusions
This paper provides new evidence on how natural gas supply shocks affect firms’ pricing decisions and inflation expectations in Italy, a country heavily reliant on natural gas as its primary energy source. We identify a gas supply shock through an external IV-VAR framework and we analyze its implications for firms’ pricing using quarterly survey data from the Bank of Italy’s SIGE, spanning over two decades. Our analysis reveals three main findings. First, natural gas supply shocks emerge as a significant driver of electricity prices in Italy. An 1% increase in the spot natural gas price measured by the Dutch TTF future determines a 0.5% increase in the Italian electricity price, measured by the PUN on impact. Second, firms respond to natural gas supply innovations by adjusting both their current and expected own prices, with inflation uncertainty rising following the shock, with effects dissipating within 5-6 quarters, consistent with the transitory nature of energy price shocks documented in the literature. The increase of inflation uncertainty in response to the shock suggests that energy price disruptions affect not only the level of prices but also firms’ ability in forecasting future inflation. Third, and most importantly, we document substantial non-nonlinearities in firms’ responses that depend on the pre-existing level of inflation uncertainty. When uncertainty is high, firms successfully pass through cost increases to consumers, maintaining elevated prices throughout the adjustment period. In contrast, when uncertainty is low, the recessionary effects of the shock dominate, causing firms to reduce prices below baseline as weak demand conditions prevent cost pass-through.