### v4

# Gas supply shocks, Uncertainty and Price setting: Evidence from Italian Firms

Abstract
This paper examines how natural gas supply shocks affect Italian firms’ pricing decisions and inflation expectations using quarterly survey data from the Bank of Italy’s Survey on Inflation and Growth Expectations (SIGE). Our findings show that gas supply shocks are quickly transmitted to spot electricity prices with persistent effects. We then estimate a proxy internalizing BVAR incorporating firm-level variables from SIGE, documenting that gas supply shocks raise firms’ current and expected prices as well as inflation uncertainty. Additionally, we document heterogeneous effects across sectors. Finally, we uncover substantial nonlinearities using state-dependent local projections: under high inflation uncertainty, firms successfully pass cost increases on to consumers, sustaining elevated prices; under low uncertainty, recessionary effects dominate, leading firms to cut prices below baseline.
Keywords: Gas supply shocks, Price Setting, Inflation Uncertainty, Firms Expectations, Local Projections, Bayesian Vector Autoregression, Instrumental Variables
JEL Codes:E31, C20, D840, C11,Q41
1 Introduction
The period that follows the Covid-19 pandemic has seen an unprecedented rise in inflation in European economies. The post pandemic recovery has been characterized by strong political tensions, dominated by events such as Russian invasion of Ukraine in the first part of 2022 which led to massive disruptions in natural gas supplies to European countries, raising the attention on inflationary role of supply shocks affecting this important commodity. Due to its limited domestic energy resources, Italy is a major importer and remains highly dependent on natural gas as a primary source of power, making it particularly sensitive to the supply disruptions that followed this dramatic geopolitical turning point. Italian firms have been particularly exposed to the unprecedented increase in the price of natural gas, which had a substantial impact on the price of electricity in the country. We show that, unsurprisingly, the post-pandemic recovery period has been characterized by an unprecedented increase in firms’ uncertainty about the future level of prices.
Unlike the extensive literature on oil supply shocks, research examining the role of natural gas supply disruptions has emerged only recently. In this paper, we aim at identifying a shock to the supply of natural gas and at evaluating the impact of this shock on firms price setting behavior. We identify the shock following an external IV-VAR approach following kanzig2021macroeconomic. We construct the instrument by exploiting episodes that carried likely exogenous news about supply disruptions affecting Europe, and in particular Italy. Our identification strategy is hence similar to the one followed by alessandri2025natural and by colombo2025understanding. Then, we evaluate the effect of this shock on firms’ pricing behavior and expectations by simulating impulse response functions from a recursive proxy internalizing VAR, using the shock identified in the first VAR as an internal instrument. We focus on firms belonging to the manufacturing and services sector, documenting important heterogeneity in the pass-trough of the shock to firms’ prices. Additionally, we evaluate the presence of nonlinearities in the pass-trough by using state dependent local projections framework (ramey2018government; F2021).
Related literature.
We draw from a vast literature about macroeconomic outcomes of energy related shocks. A large number of papers study the macroeconomic effects of oil shocks (hamilton1983oil; kilian2009not; caldara2019oil; conflitti2019oil; kilian2022oil). Our work is closely related to the recent papers which study the effect of gas supply shocks, focusing on the recent inflationary surge in European economies. alessandri2025natural identify a natural gas supply shock using daily news on European gas market as an instrument. colombo2025understanding use a similar approach for identifying gas supply shocks and extend it by identifying gas demand shocks for the US and for the Euro Area by exploiting anomalies in temperature variations. boeck2025natural identify a gas price shock for the Euro Area by using a combination of sign and zero restrictions, while lopez2025pass use sign and narrative restriction. Both of them estimate the pass through of gas price shocks to inflation. adolfsen2024gas use a BVAR with sign and narrative restrictions to identify shocks driving the natural gas market in the EU. They document that the pass through is heterogeneous depending on type of gas market related shocks hitting the economy. guntner2024sudden study the effect of recent supply disruptions in natural gas supply in the German market. casoli2024energy study the interaction between oil/gas shocks and their effects on inflation in the Euro area. With respect to those authors, we focus on Italy and on firm level expectations and price setting behavior.
Secondly, our work relates to the literature on uncertainty bloom2007uncertainty; Bloom2009, especially with regard to that related to measuring uncertainty in agents’ expectations (binder2017measuring; Jurado2015; RossiTatevik; Manski2018) and to quantifying the effects of uncertainty on the macro economy (Bloom2018; ascari2022non; GGCK2024; fasani2025belief).
The rest of the paper is organized as follows. In section 2 we describe the main features in the Survey of Inflation and Growth Expectation (SIGE), which constitute our main data source of information for firm level micro data. We then describe the aggregate statistics obtained from the survey and that we use for conducting our empirical analysis, which is described in section 3. Section 4 concludes.
2 Data: Survey on Inflation and Growth Expectations
This section illustrates the features of the Italian Survey on Inflation and Growth expectations (SIGE) [^1] https://www.bancaditalia.it/pubblicazioni/indagine-inflazione/, which constitutes our primary data source for firm expectations and price-setting decisions. As one of the longest-running firm-level expectation surveys in a G7 country, SIGE provides a rich source of information on both expectations formation and pricing behavior.
The survey is conducted by the Bank of Italy at a quarterly frequency starting from the end of 1999. It is a rotating panel that collects different kinds of information from Italian firms with at least 50 employees. The survey is stratified by (i) number of employees, (ii) geographical area, and (iii) sector. Among others, firms are asked to provide a point estimate for their year on year inflation expectations, and the expected change in their own prices over the course of the year. Furthermore, firms report the average change in their own realized price over the year, allowing a comparison between expectations and outcomes. Additionally, firms respond to a range of categorical question related to what are the main factors that will affect their own prices over the course of the next year. In this case, responses are on a scale giving information about both direction (downward or upward pressure) and intensity (ranging from strong to modest). In the following paragraphs, we describe the aggregate information that we extract from the survey in order to be used in our analysis.
Diffusion indexes
To determine which factors are most likely to impact firms’ price-setting behavior, we examine their responses to the categorical survey questions. Firms are asked about both the direction (positive, negative) and the intensity (strong, medium, modest) with which a specific factor is likely to affect their prices in the next year. The factors that are considered are the prices of raw materials and intermediate inputs, their inflation expectations for the next year, the prices of other competing firms, the trend in labor cost and aggregate demand and changes in the situation related to financing conditions. By using these responses, we build diffusion indexes using the approach of PintoSarteSharp2020. Those indexes, shown in Figure 1 together with 95% confidence bands, capture the average perceived intensity and direction of each factor across firms and are informative about movements in the distribution of firms’ responses. [^2] In order to build the indexes, each response is weighted by its intensity on a -3 to +3 scale, where higher absolute values indicate stronger perceived pressure.
Figure 1: Diffusion indexes on factors affecting firms’ future prices. The figure reports diffusion indexes computed on the categorical questions in the SIGE survey about the factors that can affect a firm’s prices in the next year, together with 95% confidence bands. Source: Survey of Inflation and Growth Expectation. Source: Bank of Italy, Survey on inflation and growth expectations. Sample: 1999:Q4–2025:Q2.
The evidence points to a clear cost-push narrative during the European energy crisis. The index for raw-material prices displays a consistently positive contribution over the entire sample, but it surges during the energy crisis, almost reaching the upper bound of the scale. This indicates an almost universal perception among firms that rising input costs were exerting strong upward pressure on their selling prices. Inflation expectations became the second most relevant upward pricing factor only around 2021, when the importance of intermediate-input costs also intensified. This pattern suggests that as firms observed widespread price increases, expectations began to feed directly into their own pricing decisions. Given the prominence of input costs—particularly energy-related raw materials—as key drivers of inflation during the post pandemic recovery, this motivates our focus on examining the effects of natural gas supply shocks on firm-level outcomes.
Firms expectations and price setting
In this paragraph, we report firms’ one-year-ahead inflation expectations, their expected price changes over the same horizon, and their actual price changes during the current year. Starting from the third quarter of 2012, an important innovation was introduced in the survey. Prior to this change, all participating firms were informed about the current level of inflation, which was reported in the survey questionnaire. Since 2012Q3, the sample size has been increased and firms have been randomly assigned to two groups: one receiving updated information about current inflation, and the other receiving no such information. Due to the length of the sample size, the aggregate statistics we report are those related to firms which have been updated about the current level of prices. It is then not surprising that the average annual inflation expectations reported in Figure 2 is closely tied to realized inflation. [^3] This stress the importance of informational frictions and inattention as key sources of bias in inflation expectations.
Figure 2: Average firms’ annual inflation expectations. The figure reports the first moment of firms’ year-on-year inflation expectations, as well as the realized HICP. Source: Survey of Inflation and Growth Expectation. Source: Bank of Italy, Survey on inflation and growth expectations, EUROSTAT. Sample: 1999:Q4–2025:Q2.
However, it is possible to observe a detachment from actual realizations during periods characterized by huge levels of economic distress. During the Lehman crisis of 2008 and the sovereign debt crisis firms tend to over predict inflation, while in the aftermath of the COVID pandemic they massively under predicted it.
Figure 3: Average change in firms’ annual expected and realized prices. The figure reports the first moment of firms’ year-on-year expectations on their own prices, as well as the annual realized change in the price they charge and realized HICP. Source: Bank of Italy, Survey on inflation and growth expectations, EUROSTAT. Sample: 1999:Q4–2025:Q2.
Figure 3 shows the average firms’ expectations for their own prices in the upcoming year, as well as the annual change in the price they charge. In general, these measures are lower than actual inflation. Both of them increase after 2020. Expected price changes peaked in 2022, in correspondence of Russian invasion of Ukraine and the start of the European energy crisis, while realized prices adjusted more sluggishly. Notably, both expected and realized prices remained at an higher level with respect to realized HICP.
(a) Price changes
(b) Expected price changes
Figure 4: Price realizations and expectations across sectors. The figure plots average price changes over the last 12 months (left) and the average expected price changes over the next 12 months (right) for Italian firms in the manufacturing (red dotted line) and services sectors (lightblue dashed line), compared with the overall series (blue solid line). Source: Bank of Italy, Survey on Inflation and Growth Expectations. Sample: 1999:Q4–2025:Q2.
In Figure 4, we compare price expectations and realizations for firms in the manufacturing and services sectors. The figures reveal a more volatile and reactive pricing behavior in manufacturing, whereas services adjust more gradually and with a slower pass-through of shocks. Until 2020, expected and realized price changes evolve similarly across sectors, following comparable cyclical patterns. The divergence emerges with the energy crisis: manufacturing firms respond much more forcefully, with both expected and realized prices rising sharply and reaching substantially higher levels than in services. This suggests that manufacturing incorporates shocks more rapidly and with greater intensity. After 2022, the adjustment in manufacturing is also more abrupt, while the correction in services is smoother and more persistent. In the post-crisis period, realized price changes in services remain comparatively higher, indicating greater rigidity than in manufacturing.
Price setting margins
We exploit firms’ answers on how much their prices changed over the previous 12 months to construct proxies of the extensive and intensive margins of price adjustment. The extensive margin is defined as the share of firms reporting a non–negligible year-on-year change in prices, which provides a survey–based measure of the incidence of price adjustment. The intensive margin is the average reported year-on-year price change among firms that adjusted. These indicators do not track the precise timing or the number of price changes within the year, but they summarize whether firms adjusted at least once over the past 12 months and by how much, and therefore offer meaningful proxies for the underlying price-setting margins.
Let $\Delta p_{i,t}^{yoy}$ denote the year-on-year percentage change in the price reported by firm $i$ in quarter $t$. We define our proxy of the extensive margin $F_{t}$ as the share of firms reporting a non-negligible price change:
$F_{t}=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\mathbf{1}\!\left(|\Delta p_{i,t}^{yoy}|\geq\varepsilon\right)$
where $\varepsilon$ is a small threshold used to filter out reporting noise. The intensive margin $I_{t}$ is instead defined as the average reported year-on-year price change among adjusting firms:
$I_{t}=\mathbb{E}\!\left[\Delta p_{i,t}^{yoy}\,\middle|\,|\Delta p_{i,t}^{yoy}|\geq\varepsilon\right].$
Under this definition, the unconditional mean of firms’ year-on-year price changes satisfies:
$\mathbb{E}[\Delta p^{yoy}_{i,t}]=F_{t}\,I_{t},$
The series, represented by the solid blue lines in Figure 5, exhibit cyclical patterns that are consistent with the micro-pricing evidence and provide a characterization of the price adjustment behavior of Italian firms over time. For both of them, we observe positive comovements with realized inflation. Before 2008, roughly 80% of firms in the sample adjusted their prices at least once per year, and the magnitude of these adjustments remained moderate. In 2008, both margins collapse: fewer firms change their prices, and those that do implement substantial price cuts, marking a sharp and broad-based disinflationary episode aligned with the onset of the global financial crisis. Following this break, the period from 2011 to 2019 is characterized by persistently muted pricing behavior. The extensive margin settles at historically low levels, while the intensive margin hovers around zero, indicating that price adjustments—when they occur—are infrequent, small, and lack a clear direction. This pattern reflects a prolonged environment of weak inflation and subdued demand. After 2020, the dynamics shift abruptly. The extensive margin rises markedly, signaling that a much larger share of firms resumes active price adjustment, while the intensive margin reaches unprecedented highs, pointing to large and predominantly upward price changes. These developments reveal a clear regime shift towards strong and widespread price increases driven by the post-pandemic recovery, the surge in energy costs, and the European gas supply shock.
(a) Extensive margin
(b) Intensive margin
Figure 5: SIGE intensive and extensive margins The figure plots the proxies for the extensive (left) and extensive (right) margins from the SIGE survey responses about the reported price changes in the last 12 months for Italian firms. Source: Bank of Italy, Survey on inflation and growth expectations. Sample: 1999:Q4–2025:Q2.
Computing the price-setting margins by differentiating across sectors reveals clear patterns. First, compared with the services sector, a slightly higher share of manufacturing firms adjust their prices over the sample. The gap between the two groups widens from 2013 until the onset of the pandemic. Manufacturing firms also start increasing their prices more frequently after 2020. In 2022, almost 90% of manufacturing firms adjusted their prices. By contrast, in the services sector the increase in the intensive margin is more gradual and contained, although still substantial.
The intensive margin is similar for both sectors for most of the sample, but after 2020 clear differences emerge: manufacturing firms adjust their prices more aggressively until 2023, after which their intensive margin decreases sharply. In contrast, price adjustments by firms in the services sector remain more contained overall, but the size of their adjustments surpasses that of manufacturing firms after 2023.
Inflation uncertainty
To measure firms’ inflation uncertainty, we adopt the approach proposed by binder2017measuring, which exploits the well-documented tendency of survey respondents to provide round-number forecasts when they face greater uncertainty. This method yields an index that captures the proportion of likely uncertain respondents in each survey wave. Differently from her case, our respondents can provide decimal answers. Therefore, we need to adapt her framework to our setting by adjusting the likelihood function accordingly. We assume that there are two type of respondents. TypehType-h respondents are characterized by an high degree of uncertainty in their subjective probability distribution, while typeltype-l respondents by a low degree of uncertainty. TypehType-h respondents are assumed to give point estimates for their expected level of prices coming from the set $S_{h}=\{0,5,10,15,20,25\}$, while low uncertainty firms are allowed to answer with decimal numbers on the support $\mathbb{R}^{+}$. If an answer $R_{it}\notin S_{h}$ we can surely classify it as typeltype-l. Conversely, if $R_{it}\in S_{h}$ we need to estimate the probability that respondent $i$ is uncertain. We assume that the cross sectional density is the mixture of two distributions. For typehtype-h respondents, coming from $S_{h}$ the probability mass function is obtained by integrating the latent density over rounding bins: $\phi_{h}(s)=P(R_{i,t}=s|type-h)=\int_{s-\delta/2}^{s+\delta/2}p_{h}(x)dx$, with $j\in S_{h}$ and $\delta$ being the bin width. [^4] We set $\delta=0.1$. For typeltype-l respondents the probability density is simply $f_{l}(x)=p_{l}(y)$ with $y\in\mathbb{R}^{+}$ and $\phi_{l}(s)=\int_{s-\delta/2}^{s+\delta/2}p_{l}(x)dx$ for rounded observations. The cross sectional log likelihood will then be given by:
$l(\theta_{t})=\sum_{s=S_{h}}n_{h}(s)log[(\lambda_{t}\phi_{h}(s)+(1-\lambda_{t})\phi_{l}(s)]+\sum_{i=1}^{N_{c}}log[(1-\lambda_{t})f_{l}(y_{c,i};\mu_{l},\sigma_{l})]$ (1)
where $f_{l}$ is assumed to be normal, $n_{h}$ is the number of firms which responses belong to $S_{h}$, $N_{c}$ is the number of non rounded responses and $\theta=(\phi_{h},\lambda_{t},\mu_{l},\sigma_{l})$ is a vector of parameters . The mixture weight $\lambda_{t}$ is the portion of uncertain respondents.555binder2017measuring includes non respondents and measures the proportion of uncertain households as $U_{t}=\frac{\lambda_{t}N_{t}+\gamma N_{t}^{DK}}{N_{t}+\gamma N_{t}^{DK}}$ where $N_{t}$ is the total number of respondents, $N_{t}^{DK}$ is the number of respondents which answer is ”don’t know” and $\gamma$ is a parameter which measures their weight. As in the SIGE survey the question about expected inflation is mandatory, we do not have DKtypeDK-type respondents in our sample and hence we measure the proportion of uncertain firms by simply taking $\lambda_{t}$. The vector of parameters $\theta_{t}$ is estimated via maximum likelihood.
Figure 6: Inflation Uncertainty Index. The figure reports the inflation uncertainty index for firms computed by adopting the framework of binder2017measuring. Vertical red bars highlight moments of high uncertainty. Source: Bank of Italy, Survey on inflation and growth expectations. Sample: 1999:Q4–2025:Q2.
Figure 6 reports the inflation uncertainty index between the last quarter of 1999 and the second quarter of 2025. Notably, since firms are informed about current price levels at the time of the interview, our index isolates doubts about future inflation, abstracting from any confusion about the current state of prices. [^6] A concern may be that building the uncertainty index only on firms who receive the information about the current level of prices may lead to an under measurement in uncertainty. In appendix A we show that the share of uncertain firms who do not receive information about current inflation when interviewed is similar both in size and temporal patterns to our baseline. Inflation uncertainty among informed firms is notably lower than that typically observed for consumers (binder2017measuring). During the first part of the sample, uncertainty remains low, reflecting stable price dynamics. The first major spike occurs in 2008Q3 during the Global Financial Crisis, followed by persistently high levels throughout the European sovereign debt crisis. Subsequent surges are observed around the Russian invasion of Crimea (2014Q3), Greece’s IMF default (2015Q2), and during the Italian banking crisis (2016Q3). The political crisis that led to the fall of the Conte II government in 2020Q4 coincides with a sharp rise in uncertainty, likely reflecting concerns over Italy’s access to European recovery funds amid the COVID-19 pandemic. A major increase appears in 2022Q1 with Russia’s invasion of Ukraine, driven by severe energy supply disruptions and their inflationary effects. Uncertainty rises again in 2022Q4, as Italian inflation reaches its highest level since the adoption of the euro, underscoring widespread concerns about the persistence and spread of price pressures. Figure 7 compares our inflation-uncertainty index with several series of interest.
Figure 7: SIGE Inflation Uncertainty Index and related series. The figure plots the inflation uncertainty index for firms computed by adopting the framework of binder2017measuring against the standard deviation of firms’ inflation expectations from SIGE (top-left), a measure of firms’ inflation uncertainty obtained by applying the procedure of Jurado2015 to the first moment of firms’ inflation expectations from SIGE and exploiting the high dimensional dataset provided by barigozzi2024large (top-right), the annualized HICP for Italy (center-left), the Economic Policy uncertainty index of baker2016measuring (center-right) and the volatility of realized inflation, estimated by adopting an ARIMA-GARCH model on the series of the annualized HICP for Italy. All series are standardized for comparability. Source: Bank of Italy, Survey on inflation and growth expectations, EUROSTAT, baker2016measuring, barigozzi2024large. Sample: 1999:Q4–2025:Q2.
First, we consider two alternative indicators of firms uncertainty about future inflation: the cross-sectional standard deviation of inflation expectations and an uncertainty index constructed following the methodology of Jurado2015, by exploiting the high-dimensional dataset of macroeconomic time series for the Italian economy provided by barigozzi2024large. [^7] No outlier correction is performed, and missing data are imputed using the algorithm of mccracken2016fred. These two measures provide natural benchmarks, as they capture either disagreement among firms or the volatility of the unpredictable component of macroeconomic dynamics. Our measure is closely aligned with these firm-specific indicators. Its correlation with the standard deviation of inflation expectations is 85.6%, and the correlation with the Jurado2015 style index is 70.1%. By contrast, correlation with annual inflation is 59.5%, and the correlation with the Economic Policy Uncertainty index of baker2016measuring is only 13.8%. This suggests that our measure is not simply reflecting changes in the level of inflation, nor is it driven by broader political or news-based uncertainty. Instead, it captures a distinct dimension of uncertainty that is specific to firms’ inflation expectations. Additionally, we compare our index with an estimate of realized inflation volatility estimated by using an ARIMA-GARCH model on annualized inflation for Italy. In this latter case, correlation with our index is 75%.
3 Empirical Analysis
3.1 Identification through external instrument
Instrument construction
In order to identify a supply shock to natural gas, we follow the approach of kanzig2023unequal (kanzig2021macroeconomic, kanzig2023unequal) based on the combination of narrative surprises and high frequency identification. As alessandri2025natural, we exploit news extracted from Refinitiv. Differently from them, we focus on a small number of news from Reuters about interruptions of gas supplies to Europe, and especially to Italy, which are highly likely to be exogenous to the economy and unexpected by the markets. Hence, we are closer to colombo2025understanding in the spirit. We include episodes of unanticipated interruptions caused by terrorist attacks, protests and unplanned maintenance in Libya and Algeria, as these two countries are important suppliers of natural gas for Italy. Furthermore, we include incidents like the explosion that hit the Baumgarten natural gas facility in Austria in 2017 causing the interruption of the flow of gas to the Italian natural gas hub in Tarvisio and the fire in the Freeport liquefied natural gas facility in US which reduced the supply of LNG to Europe. Additionally, for the 2022 period we include some announcements from Russia which have been perceived as a surprise from the market and which contain news about reduction in the supply of natural gas. As an example, the announcement by Putin that Russia was going to stop the deliveries of natural gas, unless the payments were made in robles in March 2022 concretely reduced the amount of purchasable commodity for countries which could not or did not want to comply with the request, and has been followed by a positive movement of the Dutch TTF (alessandri2025natural). We avoid to include the sabotage of the Nord Stream 2 pipeline since the latter never actually delivered natural gas to Europe, thus the event did not concretely change the amount of commodity supplied to European countries. For the same reason, we do not include the stop of gas supplies to Ukraine in 2014, as they did not affect deliveries to Europe (lopez2025pass). We also exclude the strikes in gas production facilities in Norway during 2022, since these episodes were motivated by the request of Norvegian workers to increase wages in response to inflationary pressures. Furthermore, to reduce the contamination of the instrument with geopolitical shocks we do not include the Russian invasion of Ukraine in February 2022 even though it has been followed by a massive surge in the Dutch TTF.
Table 1: Natural gas surprises
Event date Event description
1 January 2006 Russia stops flows to Ukraine due a payment dispute, blocking gas flows to Italy and Austria
22 February 2011 Due to protests and civil war, the Greenstream pipeline stopped delivering gas from Libya to Italy
31 January 2012 An unexpected cold snap forces Russia to reduce the supply of gas to Europe to satisfy its domestic demand
17 January 2013 Supplies from Algeria to Italy fall because of a terrorist attack in Almenas Gas Field
4 March 2013 Clashes with militias in Mellitha complex in Libya completely stops the gas flows to Italy.
30 September 2013 Protests in Libya cause the shut down of a gas pumping station, halting the gas flows to Italy
23 January 2014 Gas flows from Algeria to Italy stopped due to quality issues. Received gas had to be sent back
29 April 2015 Libya shuts down a gas field in the east due to protests
12 December 2017 A blast in Austrian hub of Baumgarten interrupts gas transit to Tarvisio hub in Italy
05 February 2022 An unplanned maintenance at Mellitah complex considerably reduces gas flows from Libya to Italy
23 March 2022 Putin asks European countries to pay gas deliveries in Rubles otherwise gas flows will stop
8 June 2022 Fire in the Freeport LNG facility in the US
14 June 2022 NS1 operate at limited capacity due to turbines stuck in Canada
19 August 2022 Gazprom announces that Nord Stream 1 three days shut down for maintenance on the remaining operational turbine
31 August 2022 Flows of gas from NS1 shut down permanently
2 October 2022 Russia halts gas deliveries to Italy trough the TAG pipeline
17 November 2022 Russia permanently suspends gas flows via Austria
  • Notes: The table reports the natural gas surprises we use to construct the instrument.
We end up isolating 17 surprises which are reported in Table 1. We construct the daily instrument as the log-difference of the Dutch TTF day-ahead future between the day of the event and the last trading day preceding it. Following the argument in Kanzig2020, if the risk premium remains constant within this narrow daily window, the resulting price movement can be interpreted as a news-induced revision in the expected price of natural gas. Then, we aggregate the daily instrument at the monthly level by taking the sum within the month.
IV-VAR specification
We specify a monthly VAR including six variables. We take the day ahead Dutch TTF as a measure of the spot price of natural gas. We make this price real by deflating it by a measure of the Italian HICP we built by excluding the price of natural gas. Then, in the spirit of adolfsen2024gas, we include the available gas quantities, gas inventories and industrial production for Italy. We also control for the world industrial production and for the real ”Prezzo Unico Nazionale” (PUN) as a measure of the Italian spot electricity price. [^8] Data about gas quantities and gas inventories are provided by the Italian Ministry of Environment and Energy Security. Gas quantities are defined as national production, plus imports minus exports. Gas inventories are given by the cumulative sum of the variation in inventories, given a base value of 7000 millions of standard cubic meters. Both gas quantities and gas inventories are seasonally adjusted by using the X-13 ARIMA seasonal adjustment method. Data about the PUN are provided by Gestore Mercati Energetici (GME). World industrial production is provided by BH2019. Monitoring the response of the PUN to a gas supply shock is informative because, in a uniform-price electricity market with gas-fired units often setting the marginal cost, like the Italian one, changes in the TTF day-ahead price feed into generators’ bids and hence the Italian spot price with short lags; the PUN therefore reflects downstream pass-through from gas to electricity price. Controlling for world industrial production purges the gas-price equation of global demand pressure, sharpening identification of a gas supply shock. The VAR is estimated in log levels with a constant from 2004M1 until 2025M6. The lag order is set at 12. We assess instrument relevance by regressing the VAR residuals of the TTF price on the instrument and computing the corresponding F-statistic, following olea2021inference. An F-value above 10 indicates a strong instrument. Despite the limited number of events in our narrative surprise series, the instrument is strong: in our case the F-statistic is 22.08.
Figure 8: Gas supply shock. The figure shows the natural gas supply shock obtained from the IV-VAR (left) and the corresponding correlogram (right). Sample: 2005:M4–2025:M6.
We also conduct the F-test allowing for heteroskedasticity, which delivers an F-statistic of 17.66.999The instrument explains around 8% of the TTF residual. Figure 8 reports the structural shock extracted from the IV-VAR and the corresponding correlogram.
Shock exogeneity
We test the exogeneity of our shock with respect to oil and energy-related shocks identified in the literature using a Wald test of joint significance. Specifically, we regress our shock measure on the contemporaneous and 12 lagged values of each structural innovation considered, as well as on 12 of its own lags. We consider the oil price news shock of Kanzig2020, the carbon price shock of kanzig2023unequal, the oil demand shocks of caldara2019oil and BH2019, the oil supply,inventories and demand shock of BH2019, the oil price expectation shock of Baumeister2023, and the oil supply shocks of caldara2019oil, kilian2009not and kilian2008exogenous. Results are reported in Table 2. The Wald tests do not reject the null hypotesis of no contamination for any proxy, supporting exogeneity of our shock.
Table 2: Wald test results.
Innovation Sample Wald statistic P-value
Kanzig2020 Oil price news 2005M4 - 2024M6 16.9221 0.1525
kanzig2023unequal Carbon price 2005M4 - 2019M12 13.2596 0.3505
caldara2019oil Oil demand 2005M4 - 2015M11 10.4933 0.5728
caldara2019oil Oil supply 2005M4 - 2015M11 14.1466 0.2914
BH2019 Oil supply 2005M4 - 2025M5 9.2402 0.6823
BH2019 Oil consumption demand 2005M4 - 2025M5 12.3471 0.4182
BH2019 Oil inventories 2005M4 - 2025M5 20.1281 0.0647
Baumeister2023 Oil price expectation 2005M4 - 2023M3 7.8753 0.7947
kilian2009not Oil supply 2005M4 - 2009M12 4.1983 0.9796
kilian2008exogenous Oil Supply 2005M4 - 2016M12 11.2073 0.5112
  • Notes: The table reports Wald test statistics for the joint significance of coefficients. The null hypothesis is that all coefficients are zero, indicating no contamination.
Gas supply shock transmission
Figure 9 shows the impulse response functions from the VAR estimated using the external instrument approach. As in kanzig2021macroeconomic, we can interpret the impulse response functions as elasticities, since the model is estimated in log levels. We normalize the shock to increase the gas price, measured by the TTF, by 1%. We observe that price variables immediately react to the shock. The TTF responds positively, and the response reverts to 0 in about one year. The increase in the price of natural gas is then transmitted to the national electricity price, measured by the PUN, which increases strongly, peaking within a few months and then slowly reverting. This confirms the pass-through from gas to electricity: gas-fired plants set the marginal price, so the electricity market mirrors gas dynamics with a short lag. About half of the increase in the gas price is transmitted to the electricity price in the second month after the shock. After a few months, gas inventories decrease. Higher gas prices coincide with withdrawals from storage before restocking once prices ease. Gas quantities initially do not respond to the shock on impact. After 20 months, we observe a small drop in gas quantities followed by a gradual recovery. Industrial production is not affected by the shock, consistent with its transitory nature. To verify the robustness of our results, we estimate the model including the measure of geopolitical risk by caldara2022measuring. [^10] Adding this additional variable to the model does not affect the strength of our instrument. The F statistic is 18.68 and the heteroskedasticity robust F statistic is instead 17.58. Impulse response functions for the model that includes the measure of geopolitical risk are reported in appendix B. Our baseline results are not affected by the inclusion in the model of this additional variable. This evidence suggests that our instrument is not merely capturing a generic geopolitical risk shock, but predominantly identifies shocks to the physical supply of natural gas relevant for Italy.
Figure 9: Impulse response functions of gas market variables, electricity prices and industrial production to a natural gas supply shock. The figure reports the impulse response functions of natural gas price, gas inventories, gas quantity, industrial production, PUN and world industrial production to a natural gas supply shock. Shaded areas represent 68% and 90% confidence intervals, computed using a moving block bootstrap procedure. Source: Ministero dell’ambiente e della sicurezza energetica, Gestore Mercati Energetici, ISTAT, Refinitiv Workspace, BH2019. Sample: 2004:M1–2025:M6.
To obtain a deeper understanding of how the shock propagates through the economy, we follow kanzig2021macroeconomic and re-estimate the external-proxy VAR by replacing the electricity price with a set of alternative variables, one at a time. We begin with producer price indices, distinguishing among capital, energy, and intermediate goods. Figure 10 presents the results.
Energy producer prices display a clear and statistically significant hump-shaped response: the shock raises energy input costs by roughly 0.2 – 0.4 percent over the first ten months, after which the effect gradually dissipates. Compared with the spot electricity price, the producer price index for energy exhibits a smoother and more persistent hump-shaped adjustment, consistent with the slower pass-through of energy costs through regulated tariffs, contracts and distribution margins. The response of capital-goods producer prices is modest and highly persistent, but only marginally significant. Finally, the producer price index for intermediate goods shows no statistically meaningful reaction to the shock.
Refer to captionRefer to captionRefer to caption
Figure 10: Impulse response functions of producer price indexes to a natural gas supply shock. The figure reports the impulse response functions of energy, capital goods and intermediate goods producer price indexes to a natural gas supply shock. Shaded areas represent 68% and 90% confidence intervals, computed using a moving block bootstrap procedure. Source: ISTAT.
Then, we consider a set of fuel prices provided by the Italian Ministry of Environment and Energy Security. [^11] The prices are considered net of excise and vat, and have been deflated using the italian HICP. The full set of fuel prices is available at https://sisen.mase.gov.it/dgsaie/prezzi-mensili-carburanti. All the responses for fuel prices appear to be positive, but most of them are barely significant. The only fuel that shows a statistically significant reaction is LPG. Its impulse response features a clear and persistent increase. This result is consistent with LPG being partially linked to natural gas markets and therefore more directly exposed to gas-related supply disruptions. However, this fuel has a limited importance for firms. For industrial fuels, neither light nor dense fuel oil exhibits a statistically significant reaction. Light fuel oil shows a positive median hump, but confidence bands remain wide and always include zero. Overall, the evidence indicates that LPG is the only fuel price responding significantly to gas supply shocks, while all other fuels—whether retail or industrial—show no statistically robust adjustment. This confirms that the transmission of gas supply shocks through fuel markets is generally limited in Italy and does not represent a major channel of cost pass-through for firms.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 11: Impulse response functions of real fuel prices to a natural gas supply shock. The figure reports the impulse response functions of the real price of gasoline, diesel fuel, heating diesel, LPG, dense and light fuel oil to a natural gas supply shock. Shaded areas represent 68% and 90% confidence intervals, computed using a moving block bootstrap procedure. Source: Ministry of Environment and Energy Security.
To assess whether natural gas supply disruptions propagate through financial channels, we examine equity prices, sovereign yields, short-term interest rates, and a composite indicator of systemic stress (CISS). Results are reported in Figure 12. Equity prices fall on impact and remain below baseline for several months. The 10-year government bond yield moves up in the first part of the horizon, while short-term interest rates show a small and delayed increase. The CISS reacts only marginally and quickly returns to baseline. Overall, the only responses that are statistically significant at the 68% confidence level are the initial drop in equity prices and, to a lesser extent, the short-run increase in sovereign yields, while the movements in short-term rates and systemic stress remain mostly insignificant. Overall, the evidence points to a narrow transmission of natural gas supply shocks through the energy cost channel, but its spillovers to financial markets are limited.
Refer to captionRefer to captionRefer to caption
Figure 12: Impulse response functions of share prices, 10-year goverment bond yield, short term interest rate and CISS to a natural gas supply shock. The figure reports the impulse response functions of share prices, 10-year goverment bond yield and interest rate to a natural gas supply shock. Shaded areas represent 68% and 90% confidence intervals, computed using a moving block bootstrap procedure. Source: OECD, EUROSTAT, ECB.
3.2 Natural gas supply shocks and firms’ behavior
Bayesian VAR
In order to evaluate how natural gas supply innovations affect the pricing decisions of firms, we estimate model shown in Section 3.1 and we extract the identified structural shock. Then, we estimate a proxy-internalizing VAR, placing the shock of interest as the first variable in the system and imposing recursivity (BRULU2022). [^12] According to PMW2021, the advantage of this approach over the equivalent SVAR-IV method of SW2018, besides its simplicity, is that the estimated impulse responses are valid even if the shock of interest is non-invertible. Apart from our estimated innovation, the model contains our index of inflation uncertainty, the average firms’ annual rate of change in their own prices and the average expected change in their prices over the next year, their inflation expectations and our proxies for the intensive and extensive margins. Firms expectations from SIGE are measured at quarterly frequency, at the final month of each quarter, hence the estimated gas price shock is aggregated at a quarterly level by taking the sum within the quarter. The model is estimated from 2005Q1 until 2025Q2.
Gas supply shock
[Uncaptioned image] Figure 13: Impulse response functions of firms’ related variables to a natural gas supply shock. The figure reports the impulse response functions of the inflation uncertainty index, firms’ expected and realized annual price change, inflation expectations, intensive and extensive margins and industrial confidence index to a natural gas supply shock. Shaded areas represent 68% confidence intervals. Source: Bank of Italy, Survey on inflation and growth expectations. Sample: 2005:Q1–2025:Q2.
Due to the limitations of our sample size, we resort to Bayesian methods to recover the impulse response functions. We adopt the hierarchical approach of giannone2015prior for the optimization of the hyperparameters. [^13] Since our variables in our BVAR do not exhibit any outlier in correspondence of the first and second quarter of 2020, we are not able to estimate the model by adopting the lenza2022estimate correction for accounting for increased Covid variability. Instead, we show that our IRFs are essentially unaffected when we estimate the model by completely removing the observations of 2020 from our sample. The lag-decay parameter of the Minnesota prior is treated as another hyperparameter and is estimated as well. As all our variables are stationary, we estimate the model by setting the prior mean on the coefficient of the first own lag and the prior mean of the sum-of-coefficients prior to 0. We set the number of lags to 4. Figure 13 reports the impulse response functions to a shock that increases the TTF by 1%. The results are consistent with a cost-push disturbance. On impact, inflation uncertainty increases sharply for about one year, suggesting that the shock not only pushes up the level of prices but also complicates firms’ forecasting. Firms’ prices exhibit a hump-shaped response, increasing persistently. Their response reverts to 0 in a little more than one year. The pass-through of input cost shocks to firms’ selling prices is however quantitatively limited. A gas supply shock that raises the real natural gas price by 1 percent translates into a peak increase of about 0.1 percentage points in firms’ annual price inflation. This is consistent with riggi2022price. One-year-ahead expectations about future prices display a shorter-lived increase: on impact, they increase by about 0.5 percentage points and revert to 0 within half a year. Firms’ one-year-ahead inflation expectations increase immediately on impact, showing a hump-shaped response. The response of the price margins indicates that the shock not only increases the proportion of firms that update their price; it also induces the price-adjusting firms to adjust more aggressively when they do update, as the intensive and extensive margins rise gradually. As a robustness, we compare the responses computed by adopting our baseline VAR with the ones obtained by adopting local projections. Results are reported in Figure 14. The impulse response functions delivered by the two models appear to be similar.
Gas supply shock
[Uncaptioned image] Figure 14: Impulse response functions of firms’ related variables to a natural gas supply shock. The figure reports the impulse response functions of the inflation uncertainty index, firms’ expected and realized annual price change, inflation expectations, intensive and extensive margins and industrial confidence index to a natural gas supply shock. Shaded areas represent 68% and 90% confidence intervals. Blue dotted lines represent the IRF computed by using our baseline VAR and the corresponding 68% confidence intervals. Source: Bank of Italy, Survey on inflation and growth expectations. Sample: 2005:Q1–2025:Q2.
Our results are robust to using the Jurado2015 measure of firms’ inflation uncertainty we estimate by using the high dimensional dataset for the Italian Economy provided by barigozzi2024large. Also in this case, the response of firms’ inflation uncertainty appear to be similar both in qualitative and quantitative terms. The Responses of this alternative model specification are reported in appendix C of the Appendix.
Sectoral analysis
riggi2022price document the fact that the firm size and geographical area do not appear to be factors affecting the size pass-trough of a shock affecting input prices. Instead, they find that the firm sector appear to be a relevant factor. In order to quantify the differences in the pass trough of a gas supply shock across sector, we repeat the estimation of the model shown in the previous paragraph by using the expected prices and inflation, the price changes over the last year and the extensive and the intensive margins for the manufacturing and the services sectors separately. [^14] As no significant differences emerge in terms of uncertainty among these two sector, we use the aggregate index of uncertainty for the total of the firms. This allows us to analyze differences in the price setting behavior of firms belonging to these two sectors conditionally to a natural gas supply shock. Figure 15 reports the estimated impulse response functions of the price setting variables for the two sectors. First, we see that for manufacturing firms the shock causes a markedly stronger response across all price-setting dimensions. Both realized and expected prices rise more sharply and reach their peak sooner compared with services firms, indicating that manufacturing firms adjust more aggressively to higher input costs. The intensive margin shows a similarly pronounced reaction, with manufacturing firms increasing the size of their price changes more strongly in the short run. The extensive margin also displays significant expansion, suggesting that a larger share of manufacturing firms revise their prices following the shock.
For services firms, the profile is different. The response of their own prices is barely significant and lower in magnitude with respect to firms from manufacturing sector and their inflation expectations respond with lower magnitude on impact before growing sluggisly and aligning with the one from manifacturing sector. In terms of price setting margins, the repsonse of the intensive margin is not significant, while the number of service sector firms updating their prices display a positive hump shaped response.
Overall, the results indicate that manufacturing firms react more promptly and forcefully to natural gas supply shocks, especially terms of the size of the adjustments, while the likelihood of adjusting prices is similar across the two sectors. These differences emerge despite the fact that the impulse response functions implied by those two alternative models do not imply any difference in terms of the response of inflation expectations to the shock. This finding is coherent with riggi2022price which find that inflation expectations are not an important determinant of firms’ price setting behavior, which instead are more influenced by the firms’ expected path of their own prices.
Figure 15: Impulse response functions of firms’ related variables to a natural gas supply shock. The figure reports the impulse response functions of the inflation uncertainty index, firms’ expected and realized annual price change, inflation expectations, intensive and extensive margins and industrial confidence index to a natural gas supply shock. Shaded areas represent 68% and 90% confidence intervals. Blue dotted lines represent the IRF computed by using our baseline VAR and the corresponding 68% confidence intervals. Source: Bank of Italy, Survey on inflation and growth expectations. Sample: 2005:Q1–2025:Q2.
State dependent local projections
adolfsen2024gas argue that the pass-trough of a natural gas supply shock in the Euro Area is characterized by a non linearity in the level of tightness in the labor market, with effects that are considerably magnified when the economy operates near full capacity. We show that the pass-through of natural gas supply shocks is also characterized by substantial non-linearities related firms’ ability to accurately forecast inflation. When inflation uncertainty increases, it becomes more difficult for firms to form an expectation about the future level of prices. This affects the way in which they respond to the shock. Under high uncertainty scenarios, the pass through is larger and firms persistently increase their prices. When instead uncertainty about future inflation is low they reduce their prices in order to preserve their market shares, anticipating the recessionary effects of the shock. We illustrate this by adopting state dependent local projections (ramey2018government). We determine states of high and low firms’ inflation uncertainty using the index constructed using the firm level micro data from the SIGE survey by adopting the approach of binder2017measuring. Following F2021 we scale the uncertainty index for the average inflation expectations of firms, in order to account for high inflationary periods, and we smooth it by applying a 4 period backward looking weighted moving average filter. We assume that the transition between states of high and low probabilities of inflation uncertainty is governed a logistic function $Z(\hat{\Delta}_{t-1})\in[0,1]$ of the lagged index, defined as:
$Z(\hat{\Delta}_{t-1})=\frac{exp\left(\eta\frac{\hat{\Delta}_{t-1}-\mu}{\sigma_{\hat{\Delta}}}\right)}{1+exp\left(\eta\frac{\hat{\Delta}_{t-1}-\mu}{\sigma_{\hat{\Delta}}}\right)}$ (2)
where $\hat{\Delta}_{t}$ is the state variable, $\eta$ is a parameter that determines the steepness of the transition, and $\mu$ and $\sigma_{\hat{\Delta}}$ represent the median and standard deviation of the state variable, respectively. Following F2021, we set $\eta=5$. [^15] Results are consistent across different values for parameter $\eta$. In order to account for the non linearities in unemployment shown by adolfsen2024gas, we do the same for the unemployment rate, determining the probabilities of having a tight labor market.
Figure 16: Inflation uncertainty and labor market states. Top panel: estimated probability of high inflation uncertainty following binder2017measuring, computed via smooth transition logistic function (shaded areas). Bottom panel: estimated probability of tight labor market based on unemployment, computed via smooth transition logistic function (shaded areas). Blue lines represent the underlying state variables. Source: Bank of Italy, Survey on inflation and growth expectations, EUROSTAT. Sample: 2007:Q1–2025:Q2.
The transition probabilities are reported in Figure 16, together with the smoothed and scaled index of inflation uncertainty and the unemployment rate. The probability of being in the high uncertainty state increases in the aftermath of the GFC and remains high during the sovereign debt crisis. It falls to 0 after 2019, and then it surges again in the aftermarth of the Covid-19 episode and during the European energy crisis, remaining high until the end of our sample. Instead, the probabilities of having a tight labor market are high until 2011. Then they remain at 0 until 2021, with the exception of a single episode, coinciding with the onset of the COVID pandemic in 2020. We then estimate the following set of state dependent local projections:
$x_{t+h}=$ $\bigg[\tilde{\beta}^{Z}_{x}(h)w^{gas}_{t}+\textit{controls}(x,Z)\bigg]Z(\hat{\Delta}_{t-1})$ $+\bigg[\tilde{\beta}^{\bar{Z}}_{x}(h)w^{gas}_{t}+\textit{controls}(x,\bar{Z})\bigg]\bar{Z}(\hat{\Delta}_{t-1})+\gamma_{D}D_{t}+\epsilon_{t+h}$ (3)
Where $Z(\hat{\Delta}_{t-1})$ indicate the probability of being in the state of high inflation uncertainty, $\bar{Z}\equiv 1-Z(\hat{\Delta}_{t-1})$, $\textit{controls}(x,\bar{S})$ is a set of control variables including the interaction between $w^{gas}_{t}$ and the probability of being in the tight labor market state $S(\hat{\Delta}_{t-1})$ and lags of the shock and of the dependent variable $x_{t}$. [^16] In order to preserve degrees of freedom, we set the number of lags by minimizing the AIC at horizon $h=0$, allowing for a maximum of four lags. $D_{t}$ is a dummy variable that assumes value of 1 in correspondence of 2020Q2 and then decays exponentially over the subsequent quarters. The estimated impulse response function are reported in Figure 17.
Figure 17: Firms annual price change state-dependent response to a gas supply shock. The figure shows the state dependent response of firms’ prices to a gas supply shock. Estimation follows the lag-augmentation method of MOPM2021. The number of lags of the dependent variable, the control variables and the shock in each state are determined by minimizing the AIC at horizon $h=0$. Shaded bands represent 68% and 90% confidence intervals based on Newey–West standard errors. Source: Bank of Italy, Survey on inflation and growth expectations, EUROSTAT. Sample: 2007Q1–2025Q2.
When inflation uncertainty is high, firms respond to a natural gas supply shock by persistently raising their prices. This behavior reflects precautionary pricing motives: firms face higher perceived risks about future input costs and the duration of the energy shock, and therefore widen their mark-ups to protect margins against potential future cost increases. The widespread and salient nature of energy costs also facilitates price coordination across firms, reducing competitive pressure and allowing a stronger pass-through of cost shocks. Under low uncertainty instead, firms’ prices fall in response to the shock to the gas supply. The negative effect of the shock on aggregate demand is clear. Firms anticipate the recession and do not increase their prices in order to preserve their market share.
4 Conclusions
This paper provides new evidence on how natural gas supply shocks affect firms’ pricing decisions and inflation expectations in Italy, a country heavily reliant on natural gas as its primary energy source. We identify a gas supply shock through an external IV-VAR framework and we analyze its implications for firms’ pricing using quarterly survey data from the Bank of Italy’s SIGE, spanning over two decades. Our analysis reveals three main findings. First, natural gas supply shocks emerge as a significant driver of electricity prices in Italy. An 1% increase in the spot natural gas price measured by the Dutch TTF future determines a 0.5% increase in the Italian electricity price, measured by the PUN. Second, firms respond to natural gas supply innovations by adjusting both their current and expected own prices, with inflation uncertainty rising following the shock. These effects take about one year to dissipate fully, consistently with the transitory nature of energy price shocks documented in the literature. Hence, the increase of inflation uncertainty in response to the shock suggests that energy price disruptions affect not only the level of prices but also firms’ ability in forecasting future inflation. Additionally, firms respond to the shock adjusting their prices more aggressively. Additionally, substantial differences emerge across sectors, with adjusting their prices more aggressively with respect to services sector firms. Third, we document substantial nonlinearities in firms’ responses that depend on the pre-existing level of inflation uncertainty. When uncertainty is high, firms successfully pass through cost increases to consumers, maintaining elevated prices throughout the adjustment period. In contrast, when uncertainty is low, the recessionary effects of the shock dominate, causing firms to reduce prices below baseline as weakened demand conditions prevent cost pass-through.
[…]
Appendix A Uncertainty index
Figure A.1 plots the inflation-uncertainty index, constructed following binder2017measuring, separately for three groups of firms: (i) those that were informed about current inflation at the time of the interview; (ii) those that received no information about current inflation; (iii) those that were informed only about Euro Area inflation.
Because SIGE started interviewing firms without providing current-inflation information only in the later part of the sample, the uncertainty series for groups (ii) and (iii) are substantially shorter. Nonetheless, uncertainty in these groups exhibits broadly similar dynamics to group (i). The share of firms classified as uncertain in case (ii) is about five percentage points higher on average, but this gap narrows sharply during the post-pandemic recovery. This pattern indicates that the surge in inflation uncertainty in that period is not primarily driven by informational frictions.
Figure A.1: Inflation Uncertainty Index. The figure reports the inflation uncertainty index for firms computed by adopting the framework of binder2017measuring. Vertical red bars highlight moments of high uncertainty. Source: Bank of Italy, Survey on inflation and growth expectations. Sample: 1999:Q4–2025:Q2.
Appendix B IV-VAR robustness
B.1 Controlling for geopolitical risk
Figure B.2 shows the impulse response functions obtained by augmenting our energy-sector VAR with the geopolitical risk index of caldara2022measuring and re-estimating the model using the same external instrument. The responses are virtually unchanged relative to the baseline specification. This result indicates that the instrument is not simply capturing broad geopolitical risk shocks and supports its interpretation as identifying disruptions to the physical supply of natural gas.
Gas supply shock
[Uncaptioned image] Figure B.2: Impulse response functions of energy market variables to a natural gas supply shock. The figure reports the impulse response functions of natural gas price, gas inventories, gas quantity, industrial production, PUN and world industrial production to a natural gas supply shock. Shaded areas represent 68% and 90% confidence intervals, computed using a moving block bootstrap procedure. Source: Ministero dell’ambiente e della sicurezza energetica, Gestore Mercati Energetici, ISTAT, Refinitiv Workspace, BH2019, caldara2022measuring. Sample: 2004:M1–2025:M6.
Appendix C BVAR robustness
C.1 Removing 2020 from the sample
Gas supply shock
[Uncaptioned image] Figure C.3: Impulse response functions of firms’ related variables to a natural gas supply shock. The figure reports the impulse response functions of the inflation uncertainty index, firms’ expected and realized annual price change, inflation expectations, intensive and extensive margins and industrial confidence index to a natural gas supply shock. Responses are computed by removing observations corresponding to year 2020 from the sample. Shaded areas represent 68% confidence intervals. Blue dotted lines represent the IRF computed by using our baseline VAR and the corresponding 68% confidence intervals. Source: Bank of Italy, Survey on inflation and growth expectations. Sample: 2005:Q1–2025:Q2.
C.2 Jurado measure of firms’ inflation uncertainty
Gas supply shock
[Uncaptioned image] Figure C.4: Impulse response functions of firms’ related variables to a natural gas supply shock. The figure reports the impulse response functions of the inflation uncertainty index, computed with the methodology of Jurado2015, firms’ expected and realized annual price change, inflation expectations, intensive and extensive margins and industrial confidence index to a natural gas supply shock. Shaded areas represent 68% and 90% confidence intervals. Blue dotted lines represent the IRF computed by using our baseline VAR and the corresponding 68% confidence intervals. Source: Bank of Italy, Survey on inflation and growth expectations, barigozzi2024large. Sample: 2005:Q1–2025:Q2.